### v1

# RAG-ConfusionQA: A Benchmark for Evaluating LLMs on Confusing Questions

Abstract
Conversational AI agents use Retrieval Augmented Generation (RAG) to provide verifiable document-grounded responses to user inquiries. However, many natural questions do not have good answers: about 25% contain false assumptions Yu et al. (2023), and over 50% are ambiguous Min et al. (2020). RAG agents need high-quality data to improve their responses to confusing questions. This paper presents a novel synthetic data generation method to efficiently create a diverse set of context-grounded confusing questions from a given document corpus. We conduct an empirical comparative evaluation of several large language models as RAG agents to measure the accuracy of confusion detection and appropriate response generation. We contribute a benchmark dataset to the public domain.
RAG-ConfusionQA: A Benchmark for Evaluating LLMs on Confusing Questions
Zhiyuan Peng Santa Clara University Santa Clara, CA, USA zpeng@scu.edu                        Jinming Nian Santa Clara University Santa Clara, CA, USA jnian@scu.edu                        Alexandre Evfimievski aevfim@outlook.com                        Yi Fang Santa Clara University Santa Clara, CA, USA yfang@scu.edu
1 Introduction
Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents Lewis et al. (2020); Guu et al. (2020); Gao et al. (2024). The RAG technology enables a large language model (LLM) to generate responses to users’ inquiries with information from a curated knowledge base stored as a collection of text documents. For each inquiry, a typical RAG pipeline retrieves a relevant document and uses it to generate an LLM response. Since an LLM is not a reliable source of knowledge due to its propensity to hallucinate, in business critical settings its role must be limited to language manipulation and common-sense reasoning. Every domain-specific claim in the response must be verifiable and traceable to its source document.
The performance of a RAG system is determined by evaluating its responses along a number of dimensions, including accuracy, faithfulness, completeness, answer relevance, and context relevance, as well as style metrics Es et al. (2024); Saad-Falcon et al. (2024); Liu et al. (2024); Zheng et al. (2023). Besides, the assumption that a user’s question is answerable may not hold. A well-designed RAG system should implement an exception-handling pathway that detects confusing or unanswerable questions and responds by pointing out the confusion or seeking help. Confusing questions in RAG are the subject of this work.
There is substantial prior work on using LLM to identify and respond to questions with a false premise, as well as using such questions as benchmarks for LLM evaluation Zhu et al. (2024); Yuan et al. (2024); Hu et al. (2023). Most of that work focuses on “common-sense” false premises that contradict basic knowledge learned by LLMs at pretraining. They do not consider premises that contradict a retrieved document, or appear to do so due to the lack of clarifying context such as timing. Our goal is to help LLMs respond to a variety of types of confusing questions, especially those confusing in the context of the document.
Building a “deconfusion” system involves addressing 3 challenges: creating a labeled dataset with many diverse examples of confusing questions; building a classifier to detect confusion; and, developing a proper response generator for each supported confusion type. As human content creation is expensive, we focus on LLM-based synthetic generation of confusing questions for a given set of documents. Our contributions are as follows:
  • We design, implement, and evaluate a novel synthetic data generation method to efficiently generate diverse high-quality confusing questions for a set of provided documents.
  • We discuss the types of confusing questions in RAG setting and the common situations where confusion arises.
  • We perform an empirical comparative evaluation of several LLMs as RAG agents to measure the accuracy of confusion detection and appropriate response generation.
  • We will release our code and the RAG-ConfusionQA dataset upon acceptance of the paper [^1] https://github.com/zhiyuanpeng/RAG-ConfusionQA.git.
2 Related Work
Benchmark Datasets
Most of the recent work on LLM response to confusing questions focuses on stand-alone questions to be answered based on the LLM’s pretraining knowledge, without a context document. The questions may be naturally collected from online sources Li et al. (2024); Yu et al. (2023), written by human annotators Hu et al. (2023), or synthetically generated Zhu et al. (2024); Yuan et al. (2024). Released benchmark datasets with confusing questions include KG-FPQ Zhu et al. (2024), FalseQA Hu et al. (2023), FLUB Li et al. (2024), and CREPE Yu et al. (2023). Ambiguous open-domain questions were studied in an earlier work Min et al. (2020) that released AmbigNQ, a dataset of natural ambiguous questions.
Question Generation
Zhu et al. (2024); Yuan et al. (2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of subject + relation + object using Wikipedia data, replacing the object with a similar, but wrong, object, then filling in a template to generate a false premise question. Content-grounded (non-confusing) synthetic data generation is studied in Yehudai et al. (2024); Wang et al. (2023b); Lee et al. (2023). Given the grounding passages, they generate data entries by few-shot prompting with task-specific examples, then filter to ensure data quality, faithfulness, and diversity.
False Premise Detection
by LLMs is commonly achieved with fine-tuning, chain-of-thought (CoT), and in-context learning Zhu et al. (2024); Li et al. (2024); Hu et al. (2023). Experiments show that fine-tuning outperforms the other two methods Hu et al. (2023). CoT does not bring stable improvements, while in-context learning can bring performance improvement if more examples are shown to the LLM Li et al. (2024). In Yu et al. (2023) a retrieval model finds relevant Wikipedia passages to send to a RoBERTa-based multi-passage classifier. Yuan et al. (2024) deep-dives into the LLM and finds $\sim$1of attention heads primarily responsible for getting confused by false premise questions. Constraining these attention heads during inference yields a performance increase of nearly 20%. They also observed a strong correlation between hallucination and response uncertainty metrics.
Context Based Verification
Natural questions unanswerable due to unverifiable premises were studied in Kim et al. (2021). This work is closer to RAG because it associates a document with each question: a Wikipedia page retrieved by the Google search engine. A premise is unverifiable (rather than false) if it not entailed by the document. The proposed system extracts all premises from a question, verifies them by checking entailment from each sentence of the document, then generates a response stating the unverifiable premise if found.
3 Methods
We built benchmark RAG-ConfusionQA to evaluate LLMs on the context-grounded response or answer generation step of the RAG pipeline when the questions are confusing (Section 3.1). In this step, given a standard prompt $p$ (Appendix F.6), a confusing question $q$, and a retrieved document $d$, the LLMr generates a response $r$. Ideally, $r$ should clarify the confusing part of the question, a process we refer to as defusion (or de-confusion), rather than attempting to answer the question directly. Document retrieval and ranking are out of scope for this work.
3.1 Confusing Questions
We call question $q$ confusing given document $d$ if $q$ cannot be answered faithfully and completely based on $d$ alone, without requesting — or providing — some clarification. Common confusion types in RAG-ConfusionQA (Appendix D) include:
False premise:
$q$ makes a wrong or a highly biased assumption that must be clarified. The contradiction may be with $d$ or with basic knowledge.
Not Mention:
Some in-domain information required to answer $q$ is not provided in $d$.
Ambiguous:
$q$ is missing some context or contains generic words with multiple interpretations, allowing different answers in the context of $d$.
Note that $q$ is interpreted within the context of $d$. If the name “John Smith” appears in both, it is assumed to refer to the same person unless stated otherwise. Similarly, $q$ and $d$ are assumed to refer to the same time, place, or use-case unless clarified, otherwise leading to confusion.
Algorithm 1 The guided hallucination method
1:Call LLMq to extract from $d$ a list of $n$ claims $c_{1},\ldots,c_{n}$; batching claims reduces LLMq-calls (Appendix F.1)
2:Partition $\{1,\ldots,n\}$ into disjoint subsets $S_{1},\ldots,S_{k}$, e.g. $S_{j}=\{i:i\mathop{\textrm{mod}}3=j-1\}$ for $j=1,2,3$
3:for several rounds (e.g. 3), and in each round, for all $S_{j}$ in partition $(S_{1},\ldots,S_{k})$ do
4:     In the claims list $c_{1},\ldots,c_{n}$ replace all claims $c_{i}$ where $i\in S_{j}$ by text “(missing)
5:     Provide LLMq with this modified list (but not $d$) and prompt it to recover the missing claims (Appendix F.2)
6:     In the claims list $c_{1},\ldots,c_{n}$ replace the missing claims with the imputed claims
7:end for
8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only novel claims (Appendix F.9)
9:Call LLMq to generate one short question per each novel claim, focusing on one key element of it (Appendix F.8)
10:Call LLMq to filter out any questions answerable in the context of $d$ (Appendix F.3)
3.2 Data Generation
Here we describe our method of generating confusing questions for a given set of documents. We require each document to be short enough to fit in the LLM prompt along with instructions and examples, but long enough to make at least 5 to 10 separate claims (i.e. a few paragraphs long). We prefer documents that are novel to the LLMs under evaluation, e.g. news articles published later than all pretraining cutoff dates, so that the LLMs cannot reproduce claims or answer questions correctly by pure hallucination. Thus, we collected 200 news articles published in 2024, each containing more than 150 words for each topic (Appendix E). Sentences were selected sequentially from the top until the total word count exceeded 300, using the NewscatcherBugara et al. (2020). Given a set of such documents, we generate multiple confusing questions about each document, along with some non-confusing ones for comparison.
We start by picking LLMq (GPT-4o-mini by default in this paper), the “teacher” LLM that will generate the confusing questions. We prefer, but do not require, LLMq to be larger and stronger than the LLMr under evaluation. For each document $d$, we call LLMq to generate a list of factual claims, then iteratively replace subsets of claims with hallucinated claims while maintaining semantic similarity to $d$ using guided hallucination, see Algorithm 1. We also generate some non-confusing questions for each document $d$ by calling LLMq either directly.
3.3 Evaluation and Metrics
We first assessed the quality of RAG-ConfusionQA by sampling a subset of questions and manually verifying whether they were indeed confusing or not (Section 3.4). Next, we evaluated how well various LLMr models could defuse the confusing questions in RAG-ConfusionQA. Given the large number of confusing questions, manual verification of LLMr responses is not feasible. Therefore, we proposed AutoDefuseEval using GPT-4o-mini to automatically detect whether a response successfully defuses the question. To validate the effectiveness of this defusion detection method, we compared its performance with human annotations on a sampled dataset (Section 3.4). Formally, we define our tasks as follows:
Question generation:
What is the quality of confusing and non-confusing questions generated by Algorithm 1? (Section 3.4)
Defusion detection:
How accurately can AutoDefuseEval detect if an LLMr response to a confusing question defuses the confusion? (Section 3.4)
Confusion response:
How frequently does an LLMr, when responding to a confusing question using the standard prompt, generate a response that successfully defuses the confusion? (Section 4.1)
Confusion detection:
How accurately can LLMs identify which context-grounded questions are confusing and require special handling? (Section 4.2)
We evaluate “question generation” and “defuse detection” through accuracy and confusion matrix. We only apply accuracy for “confusion response” and “confusion detection”. Algorithm A has the steps we run for each LLMr and for each $(q,d)$ pair. We run steps 1, 2, and 3 for “confusion detection” and 1, 4, 5, and 6 for “confusion response”.
Following the self-consistency method Wang et al. (2023a), we perform multiple LLM calls in steps 2 and 5 of Algorithm A, taking the majority vote to determine the final label. We also present the LLMq with few-shot exemplars of questions, responses, and explanations (Section 4.3).
Model business entm food music news politics science sport tech travel Avg Std Dev
GPT-3.5 18.90 12.39 13.55 15.48 15.15 17.13 10.81 $20.24^{\star}$ 17.12 10.01 15.08 3.21
Llama 3.2 3B 68.34 66.11 64.89 67.90 67.50 66.94 57.98 $76.37^{\star}$ 67.07 60.06 66.31 4.70
Mistral 7B v0.3 71.27 70.37 65.81 71.60 71.62 69.26 60.35 $79.00^{\star}$ 69.97 64.29 69.35 4.77
Llama 3.1 70B 74.48 72.55 70.33 75.31 71.33 71.39 65.60 $\textbf{83.33}^{\star}$ 73.17 70.38 72.79 4.34
Llama 3.1 8B 76.18 74.03 70.23 76.07 73.15 72.87 64.37 $81.45^{\star}$ 73.47 68.11 72.99 4.45
Table 1: Evaluation of LLMs’ accuracy in defusing confusing questions from the RAG-ConfuseQA-Silver dataset across diverse news topics. For each LLM, the underscored value and stared value are the minimum and maximum values, respectively, across all the topics. Bold values are the maximum values across all the LLMs.
3.4 Human Annotation
To evaluate “question generation” and “defuse detection” (Section 3.3), we recruited six computer science graduate students for manual labeling (Appendix B). As reflected in “Ground Truth Acc,” of Appendix B, Table  3, our method achieved 94.91% accuracy in generating confusing and non-confusing questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion. As shown in Appendix B, Figure 2, the confusion matrix on the left strongly agrees with the ground truth, with minimal errors (5 false negatives and 6 false positives). The defusion matrix on the right reflects a similarly high accuracy, with only one error for false negatives and false positives.
4 Experimental Results
4.1 Confusion Response
We selected several open-source LLMs (Appendix 4) to evaluate their ability to generate responses that effectively defuse confusing questions with a standard prompt (Appendix F.6). As shown in Table 1, a) GPT-3.5 performs the worst among all six LLMs but with the smallest Std Dev across all topics, indicating consistent performance. On the other hand, Llama 3.1 70B and Llama 3.1 8B achieve comparable top performances with similar Std Dev, despite their differences in model size; b) most LLMs perform the worst in the “science” topic, except for GPT-3.5, for which “science” is the second worst. This is expected, as science topics often require more specialized domain knowledge to defuse confusion effectively; c) all LLMs achieve their best performance on the “sport” topic, demonstrating a clear trend. The strong performance across LLMs on the “sport” topic might be attributed to the relative simplicity and unambiguous nature of sports-related content.
4.2 Confusion Detection
By default, we use GPT-4o-mini as our LLMq for the task of “confusion detection,” which identifies whether the generated questions are confusing. In this section, we compare GPT-4o-mini with other LLMs. As illustrated in Figure 1, GPT-4o-mini achieves the best performance, while the smallest model, Llama 3.2 3B, performs the worst.
Llama 3.2 3BMistral 7BLlama 3.1 8Bgpt-3.5Llama 3.1 70Bgpt-4o-mini606060608080808010010010010079.6379.6379.6379.6385.1985.1985.1985.1986.5786.5786.5786.5787.9687.9687.9687.9693.5293.5293.5293.5294.9194.9194.9194.91Different LLMsAccuracy (%)
Figure 1: Evaluation of instruction-finetuned LLMs’ accuracy in classifying questions as confusing or not on the RAG-ConfusionQA-Gold dataset
Module Accuracy
Base (gpt-4o-mini) 87.61
+ examples 97.35
+ self-consistency (n=3) 98.23
+ self-consistency (n=9) 98.23
Table 2: Evaluation of different prompts in the task “defusion detection” on RAG-ConfusionQA-Golden.
4.3 Ablation Study
We examined the impact of various prompts on the task of “defusion detection.” As seen in Table 2, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $n=3$ further boosted performance to 98.23%. However, increasing the number of generated sequences beyond three did not lead to additional improvements. Since RAG-ConfusionQA-Golden is a smaller dataset, using $n=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger RAG-ConfusionQA-Silver dataset. Therefore, we adopted $n=9$ in our experiments to balance accuracy and expenses (Appendix F.10).
5 Conclusion
We introduced RAG-ConfusionQA, a benchmark designed to assess how effectively an LLM defuses confusion when responding to confusing questions using a standard prompt. Through human annotation, we demonstrated that RAG-ConfusionQA achieves a 94.91% accuracy in generating confusing and non-confusing questions, while our proposed AutoDefusionEval achieves a 98.23% accuracy in detecting whether an LLM’s response defuses confusion. Additionally, our evaluation highlights that current LLMs often attempt to answer confusing questions to varying degrees. For future work, we aim to develop methods to better detect confusing questions and address the tendency of LLMs to respond to them directly instead of defusing the confusion.
6 Acknowledgement
We would like to express our sincere gratitude to Xiaoxiao Shang, Ethan Lin, Wei Mo, Zhan Shi, Xinze Ye, and Haisong Wang for their invaluable contributions to data labeling.
7 Limitations
In this paper, we only evaluated LLMs’ abilities to generate responses that effectively defuse confusing questions when prompted with a standard prompt (Appendix F.6). We chose this prompt because, ideally, LLMs should be able to address confusion with a basic, straightforward prompt. However, we have not yet assessed the performance of LLMs across different prompt variations.
Appendix A Algorithm
Algorithm 2 Evaluation steps
1:Call LLMr on $(d,q)$ with the standard RAG prompt and get its response $r$ (Appendix F.6)
2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is confusing given $d$ (Appendix F.3)
3:Aggregate $m$ predictions by majority vote
4:if $q$ is generated as confusing then
5:     Call LLMq $m$ times on $(d,q,r)$ to check and explain if $r$ defused the confusion (Appendix F.4)
6:     Aggregate $m$ predictions by majority vote
7:end if
Appendix B Human Annotation
Figure 2: Confusion matrix of confusion and defusion on RAG-ConfusionQA-Golden.
Group Name Cohen’s Kappa Annotator Acc Group’s Agree Acc Ground Truth Acc
confusion defusion confusion defusion confusion defusion confusion defusion
1   — 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
  — 87.50 94.59
2   — 0.7508 0.8187 93.06 97.14 98.41 96.77
  — 91.67 90.00
3   — 0.8615 0.8333 91.67 93.75 98.51 100.00
  — 98.61 100.00
Table 3: Agreement scores and the performance of our method on the RAG-ConfusionQA-Golden.
Following the methodology from Cochran (1977), representing approximately 20,000 synthetic questions in RAG-ConfuseQA at a 95% confidence level, with a 4% margin of error, and an anticipated accuracy of $90\pm 5\%$, required labeling at least 202 questions. We sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators. As shown in Table 3, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is confusing and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency. Two more annotators manually resolved all disagreements to establish ground truth labels. Six annotators are computer science graduate students recruited via email and incentivized with complimentary meals. Another two annotators are the authors of this paper, who are also majoring in computer science.
We used two metrics to assess our method: “Annotator Acc” and “Group’s Agree Acc.” “Annotator Acc” treats each annotator’s labels as ground truth, computing the accuracy of our method’s generated labels. “Group’s Agree Acc” uses the instances where both annotators within the same group agreed, treating this consensus as the ground truth. This method demonstrated better accuracies than “Annotator Acc,” leading us to resolve disagreements with two additional annotators, ultimately establishing the final labeled data as ground truth.
Appendix C Different LLMr
Model Name Quantization Knowledge Cut-off Date
Llama 3.2 3B Instruct Turbo FP8 2023/12
Llama 3.1 8B Instruct Turbo FP8 2023/12
Llama 3.1 70B Instruct Turbo FP8 2023/12
Mistral (7B) Instruct v0.3 FP16 2024/5
gpt-3.5-turbo   — 2021/9
Table 4: Different LLMr to be evaluated. In RAG-ConfuseQA-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cut-off date 5/22/2024.
Appendix D Confusing Types
Document:
Justin Jefferson, CeeDee Lamb, NFL Injury Statuses and Fantasy Impact for Week 3 Stephen Maturen/Getty Images
Two of the best wide receivers in the NFL are expected to be on the field for their respective Week 3 games. Justin Jefferson and CeeDee Lamb each received positive prognosis about their injury issues, which is something that can’t be said for the rest of the stars across the NFL. Christian McCaffrey is already on injured reserve, Deebo Samuel is out for a weeks and a slew of other running backs and wide receivers are dealing with ailments that could keep them out of Week 3. Below is a look at all of the significant injuries that could affect fantasy football matchups across Week 3. Justin Jefferson Off Injury Report
Justin Jefferson was taken off the Minnesota Vikings injury report on Friday. Jefferson’s status was up in the air because of a quad injury, but he practiced well enough this week that the injury is not a concern. The superstar wide out will be needed for Minnesota’s home clash with the Houston Texans, which has the potential to be a high-scoring affair. Jefferson is always the primary target in Minnesota when healthy, but he should have more targets in Week 3 because Jordan Addison and T.J. Hockenson are still out. Jefferson earned seven targets from Sam Darnold in Week 2. Only running back Aaron Jones had more than four targets against the New York Giants. Houston’s defense may give Jefferson some fits, led by cornerback Derek Stingley Jr. but it has allowed 330 receiving yards to opposing wide outs on just 20 catches through two weeks. Jefferson is an automatic start whenever he’s healthy, and his star power may be needed more on certain fantasy rosters in Week 3 depending on how many injuries affect a single roster.
Question 1: How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?
This question is confusing as it asks about "Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft" which was never discussed in the document. So this belongs to Not Mention
Question 2: How will Justin Jefferson’s struggles with hamstring issues in Week 3 affect the Vikings’ chances of winning against the Houston Texans?
This question assumes that Justin Jefferson is dealing with hamstring issues, which is not mentioned in the document. The document only refers to a quad injury that Jefferson had, which is no longer a concern as of Week 3. The document makes it clear that Jefferson is healthy and will play without any restrictions. Therefore, the question is based on a False Premise
Question 3: How will his performance impact the team’s chances this season, considering they are missing key players?
This question uses the Ambiguous pronoun “his” without specifying who it refers to, making it unclear whether the question is asking about Justin Jefferson, CeeDee Lamb, or another player mentioned in the document. Additionally, “the team” is not defined explicitly, as both the Minnesota Vikings and other teams are referenced in the document. Without clarifying which player or team the question is referring to, it is impossible to provide an accurate answer based on the document.
Appendix E Statistics of RAG-ConfusionQA
Attr Golden Silver
# of documents 45 2,000
# of non-confusing questions 103 9,052
# of confusing questions 113 10,218
average words per document 280 276
average words per non-confusing question 16 16
average words per confusing question 16 16
Table 5: Statistics of RAG-ConfusionQA. Golden is sampled from Silver and annotated by annotators (Appendix B).
Appendix F Prompt
F.1 Extract Claims Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and list {num_fact} most important facts it contains. Each fact should be
stated in a clear, standalone sentence with sufficient context to be understood
independently, avoiding undefined pronouns. Ensure that each fact is directly derived from
the document and does not include any information not mentioned within it.
Document:
    """{document}"""
{num_fact} most important facts:"
F.2 Recover Missing Claims Prompt
{mdframed}
Read the document below with a list of {num_fact} facts it contains. Note that some of the
facts are missing represented by (missing). Your task is to guess the missing facts could
have said and complete the missing facts. Each fact should be stated in a clear, standalone
sentence with sufficient context to be understood independently, avoiding undefined pronouns.
Please complete the missing facts and return all the {num_fact} facts in original order. You
must return all the {num_fact} facts.
Document:
    """{document}"""
Completed list of facts:"
F.3 Confusion Judgement Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and the question. Check if the question contains any confusing
parts, false assumptions, incorrect information, or mentions of people, locations,
organizations, or other entities that are not present in the document.
Document:
    """{document}"""
Question:
    {question}
Does the question contain any confusing parts, false assumptions, incorrect
information, or mentions of people, locations, organizations, or other entities
that are not present in the document? Please start by explaining your reasoning.
End your response with ’The answer is: Yes or No.’:
F.4 Defusion Judgement Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and answer the question based on the document.
Document:
    """{example_document}"""
Question:
    {example_question}
{defuse_answer_examples}
{Defusion Instruction}
{defuse_reasoning_examples}
Read the document and answer the question based on the document.
Document:
    """{document}"""
Question:
    {question}
{response_from_gpt3.5}
{Defusion Instruction}
F.5 Defusion Instruction
{mdframed}
The question presented is confusing and unanswerable based on the document provided.
Read the response carefully. If the response attempts to answer the question,
like answering ’Yes’ to agree the question or ’No’ to disagree the question,
you should immediately conclude with The answer is: No. even if it later
acknowledges the questions confusion. Otherwise, determine if the response
clarifies that the question cannot be answered or confusing or matches any of the
issues listed:
    False Premise: The response points out that the question is based
    on an incorrect assumption.
    Not Mentioned: The response indicates that the information needed to answer
    the question is not available or not mentioned in the document.
    Ambiguous: The response points out unclear pronouns or vague terms in the
    question that prevent a clear answer.
Begin with your reasoning and finally conclude with The answer is: Yes. or
The answer is: No.. Note that if the response attempts to answer the question
at any point, like answering ’Yes’ to agree the question or ’No’ to disagree the
question, you should immediately conclude with The answer is: No. even if it
later acknowledges the questions confusion:
F.6 RAG Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and answer the question based on the document.
Document:
    """{document}"""
Question:
    {question}
F.7 Non-confuing Question Generation Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document attentively and compile a numbered list of the top {num_q} questions
that the document directly answers. Ensure each question is clear, accurate, and devoid
of confusion, false assumptions, undefined pronouns, or misinformation. Avoid referencing
people, locations, organizations, or other entities not explicitly mentioned in the
document. Construct each question to be thought-provoking, containing between 13 to 18
words, and sufficiently detailed to avoid being overly straightforward.
Document:
    """{document}"""
Questions:
F.8 Confusing Question Generatioin Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document and review the list of hallucinated facts. For each hallucinated fact,
craft a single, specific and concise question containing 13 to 18 words that incorporates
the key element of the fact, ensuring the question is intentionally confusing.
The question should not be answerable using any information present in the document.
The question should not combine multiple queries and each question should address only
one specific aspect. If a question cannot be formulated for a particular hallucinated
fact, you may omit it.
Document:
    """{document}"""
hallucinated facts:
    {hallucinated_facts}
Questions:
F.9 Remove Claims Prompt
{mdframed}
You will be provided with a document delimited by triple quotes.
Read the document and follow user’s instructions.
Read the document below with a list of {num_true_fact} ground-truth facts it contains
and a list of {num_false_fact} hallucinated facts that are not supported by the document.
Your task is to remove any hallucinated facts that can be supported by either the
document or the {num_true_fact} ground-truth facts.
Please only return the remaining hallucinated facts, along with their original order numbers.
Document:
    """{example_document}"""
4 ground-truth facts:
    {true_facts}
{num_false_fact} hallucinated facts:
    {false_facts}
Remaining hallucinated facts:
    {remained_facts}
Read the document below with a list of {num_true_fact} ground-truth facts it contains
and a list of {num_false_fact} hallucinated facts that are not supported by the document.
Your task is to remove any hallucinated facts that can be supported by either the
document or the {num_true_fact} ground-truth facts.
Please only return the remaining hallucinated facts, along with their original order numbers.
Document:
    """{document}"""
{num_true_fact} ground-truth facts:
    {true_facts}
{num_false_fact} hallucinated facts:
    {hallucinated_facts}
Remaining hallucinated facts:
F.10 Expenses
Generating the RAG-ConfusionQA dataset using GPT-4o-mini costs approximately \$50. For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/. Excluding the dataset generation, all other experiments cost around \$60.

### v2

@@
-RAG-ConfusionQA: A Benchmark for Evaluating LLMs on Confusing Questions
+ScopeQA: A Framework for Generating Out-of-Scope Questions for RAG
 Abstract
@@
 RAG agents need high-quality data to improve their responses to confusing questions.
-This paper presents a novel synthetic data generation method to efficiently create a diverse set of context-grounded confusing questions from a given document corpus.
+This paper presents a novel guided hallucination-based method to efficiently generate a diverse set of borderline out-of-scope confusing questions for a given document corpus.
 We conduct an empirical comparative evaluation of several large language models as RAG agents to measure the accuracy of confusion detection and appropriate response generation.
 We contribute a benchmark dataset to the public domain.
-RAG-ConfusionQA: A Benchmark for Evaluating LLMs on Confusing Questions
-Zhiyuan Peng Santa Clara University Santa Clara, CA, USA zpeng@scu.edu                        Jinming Nian Santa Clara University Santa Clara, CA, USA jnian@scu.edu                        Alexandre Evfimievski aevfim@outlook.com                        Yi Fang Santa Clara University Santa Clara, CA, USA yfang@scu.edu
+ScopeQA: A Framework for Generating Out-of-Scope Questions for RAG
+Zhiyuan Pengthanks: Equal contribution.
+Santa Clara University Santa Clara, CA, USA zpeng@scu.edu                        Jinming Nian11footnotemark: 1 Santa Clara University Santa Clara, CA, USA jnian@scu.edu                        Alexandre Evfimievski aevfim@outlook.com                        Yi Fang Santa Clara University Santa Clara, CA, USA yfang@scu.edu
 1 Introduction
@@
 (2024).
-The RAG technology enables a large language model (LLM) to generate responses to users’ inquiries with information from a curated knowledge base stored as a collection of text documents.
-For each inquiry, a typical RAG pipeline retrieves a relevant document and uses it to generate an LLM response.
-Since an LLM is not a reliable source of knowledge due to its propensity to hallucinate, in business critical settings its role must be limited to language manipulation and common-sense reasoning.
-Every domain-specific claim in the response must be verifiable and traceable to its source document.
-The performance of a RAG system is determined by evaluating its responses along a number of dimensions, including accuracy, faithfulness, completeness, answer relevance, and context relevance, as well as style metrics Es et al.
+Given an inquiry, a RAG system retrieves a relevant document from a curated knowledge base and provides it to a large language model (LLM) that generates a response.
+RAG performance is evaluated on dimensions like the response accuracy, faithfulness, completeness, answer and context relevance, and style Es et al.
 (2024); Saad-Falcon et al.
@@
 (2023).
-Besides, the assumption that a user’s question is answerable may not hold.
-A well-designed RAG system should implement an exception-handling pathway that detects confusing or unanswerable questions and responds by pointing out the confusion or seeking help.
-Confusing questions in RAG are the subject of this work.
-There is substantial prior work on using LLM to identify and respond to questions with a false premise, as well as using such questions as benchmarks for LLM evaluation Zhu et al.
-(2024); Yuan et al.
-(2024); Hu et al.
-(2023).
-Most of that work focuses on “common-sense” false premises that contradict basic knowledge learned by LLMs at pretraining.
-They do not consider premises that contradict a retrieved document, or appear to do so due to the lack of clarifying context such as timing.
-Our goal is to help LLMs respond to a variety of types of confusing questions, especially those confusing in the context of the document.
-Building a “deconfusion” system involves addressing 3 challenges: creating a labeled dataset with many diverse examples of confusing questions; building a classifier to detect confusion; and, developing a proper response generator for each supported confusion type.
-As human content creation is expensive, we focus on LLM-based synthetic generation of confusing questions for a given set of documents.
+Errors may stem from a document-to-question mismatch, hallucinating parts of the response, wrong understanding of the document or the question.
+Many user questions have no direct answer: some may rely on a false premise, some may be ambiguous in what they ask, and some are out of scope for the knowledge base.
+A well-designed RAG system should detect such confusing questions and respond appropriately, either by clarifying the confusion or seeking help.
+Dataset RAG Setting Source Category Cut-off-knowledge Annotation Method
+AMBIGQA Min et al. (2020) Wiki Multi-answers Human
+CREPE Yu et al. (2023) Reddit False Premise Human
+FalseQA Hu et al. (2023)  False Premise Human
+KG-FPQ Zhu et al. (2024) Wiki False Premise LLM
+FAITH Yuan et al. (2024) Wiki False Premise LLM
+FLUB Li et al. (2024a) Baidu Tieba Cunning Texts Human
+ScopeQA News Out-of-scope Human+LLM
+Table 1: ScopeQA vs Existing Datasets
+Enabling “deconfusion” in a RAG system involves three challenges: (1) creating a diverse dataset of confusing questions, (2) training a classifier to detect confusion, and (3) developing a response generator for each confusion type.
+As shown in Table 1, much prior work focuses on false-premise questions that contradict pretraining knowledge but neglects out-of-scope queries, which lack answers in retrieved documents.
+Borderline out-of-scope questions can confuse LLMs into making up a hallucinated answer.
+This type of confusion is the focus of this paper.
+While queries in MS MARCO Nguyen et al.
+(2016) may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cut-off.
+ScopeQA fills this gap by providing a diverse benchmark for out-of-scope questions, leveraging LLM-assisted generation to minimize human annotation effort.
 Our contributions are as follows:
-  • We design, implement, and evaluate a novel synthetic data generation method to efficiently generate diverse high-quality confusing questions for a set of provided documents.
-  • We discuss the types of confusing questions in RAG setting and the common situations where confusion arises.
-  • We perform an empirical comparative evaluation of several LLMs as RAG agents to measure the accuracy of confusion detection and appropriate response generation.
-  • We will release our code and the RAG-ConfusionQA dataset upon acceptance of the paper [^1] https://github.com/zhiyuanpeng/RAG-ConfusionQA.git.
+  • We propose and implement a method to generate borderline out-of-scope questions for given documents efficiently.
+  • We validate our synthetic data by training out-of-scope question detectors.
+  • We evaluate LLMs as RAG agents for detecting confusion and generating responses.
+  • Our code and dataset are released at https://github.com/zhiyuanpeng/ScopeQA.git.
 2 Related Work
 Benchmark Datasets
-Most of the recent work on LLM response to confusing questions focuses on stand-alone questions to be answered based on the LLM’s pretraining knowledge, without a context document.
+Most of the recent work on LLM responses to confusing questions focuses on stand-alone questions answered using LLM’s parametric knowledge without a context document.
 The questions may be naturally collected from online sources Li et al.
-(2024); Yu et al.
+(2024b); Yu et al.
 (2023), written by human annotators Hu et al.
@@
 (2024).
-Released benchmark datasets with confusing questions include KG-FPQ Zhu et al.
-(2024), FalseQA Hu et al.
-(2023), FLUB Li et al.
-(2024), and CREPE Yu et al.
-(2023).
-Ambiguous open-domain questions were studied in an earlier work Min et al.
-(2020) that released AmbigNQ, a dataset of natural ambiguous questions.
+A comparison of these released benchmark datasets is shown in Table 1.
 Question Generation
@@
 (2024); Yuan et al.
-(2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of subject + relation + object using Wikipedia data, replacing the object with a similar, but wrong, object, then filling in a template to generate a false premise question.
+(2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of subject + relation + object from Wikipedia, replacing the object with a similar but incorrect one, and filling in a template.
 Content-grounded (non-confusing) synthetic data generation is studied in Yehudai et al.
@@
 Given the grounding passages, they generate data entries by few-shot prompting with task-specific examples, then filter to ensure data quality, faithfulness, and diversity.
-False Premise Detection
-by LLMs is commonly achieved with fine-tuning, chain-of-thought (CoT), and in-context learning Zhu et al.
+Mitigation
+is commonly achieved with in-context learning  Zhu et al.
 (2024); Li et al.
-(2024); Hu et al.
-(2023).
+(2024b), chain-of-thought (CoT)  Wei et al.
+(2022) and fine-tuning.
 Experiments show that fine-tuning outperforms the other two methods Hu et al.
 (2023).
-CoT does not bring stable improvements, while in-context learning can bring performance improvement if more examples are shown to the LLM Li et al.
-(2024).
-In Yu et al.
-(2023) a retrieval model finds relevant Wikipedia passages to send to a RoBERTa-based multi-passage classifier.
+CoT yields inconsistent improvements, while in-context learning improves performance with more examples Li et al.
+(2024b).
 Yuan et al.
-(2024) deep-dives into the LLM and finds $\sim$1of attention heads primarily responsible for getting confused by false premise questions.
-Constraining these attention heads during inference yields a performance increase of nearly 20%.
-They also observed a strong correlation between hallucination and response uncertainty metrics.
-Context Based Verification
-Natural questions unanswerable due to unverifiable premises were studied in Kim et al.
-(2021).
-This work is closer to RAG because it associates a document with each question: a Wikipedia page retrieved by the Google search engine.
-A premise is unverifiable (rather than false) if it not entailed by the document.
-The proposed system extracts all premises from a question, verifies them by checking entailment from each sentence of the document, then generates a response stating the unverifiable premise if found.
+(2024) identifies  1% of attention heads responsible for confusion by false-premise questions, showing a 20% performance gain by constraining these heads.
+Kim et al.
+(2021) retrieves a Wikipedia page for each question and responds with the identified unverifiable premises if the page does not entail them.
 3 Methods
-We built benchmark RAG-ConfusionQA to evaluate LLMs on the context-grounded response or answer generation step of the RAG pipeline when the questions are confusing (Section 3.1).
-In this step, given a standard prompt $p$ (Appendix F.6), a confusing question $q$, and a retrieved document $d$, the LLMr generates a response $r$.
-Ideally, $r$ should clarify the confusing part of the question, a process we refer to as defusion (or de-confusion), rather than attempting to answer the question directly.
-Document retrieval and ranking are out of scope for this work.
-3.1 Confusing Questions
-We call question $q$ confusing given document $d$ if $q$ cannot be answered faithfully and completely based on $d$ alone, without requesting — or providing — some clarification.
-Common confusion types in RAG-ConfusionQA (Appendix D) include:
-False premise:
-$q$ makes a wrong or a highly biased assumption that must be clarified. The contradiction may be with $d$ or with basic knowledge.
-Not Mention:
-Some in-domain information required to answer $q$ is not provided in $d$.
-Ambiguous:
-$q$ is missing some context or contains generic words with multiple interpretations, allowing different answers in the context of $d$.
-Note that $q$ is interpreted within the context of $d$.
-If the name “John Smith” appears in both, it is assumed to refer to the same person unless stated otherwise.
-Similarly, $q$ and $d$ are assumed to refer to the same time, place, or use-case unless clarified, otherwise leading to confusion.
+We developed the ScopeQA data generator to evaluate and mitigate LLM confusion when responding to out-of-scope questions.
+We assume that the retrieved document is the best available match for the question and focus on the response generation step of the RAG pipeline.
+In this step, given a standard prompt $p$, an out-of-scope question $q$, and a retrieved document $d$, the LLMr generates a response $r$.
+Ideally, $r$ should clarify the confusing part of the question, a process we refer to as defusion (or de-confusion), rather than attempting to answer the question directly and hallucinate.
 Algorithm 1 The guided hallucination method
 1:Call LLMq to extract from $d$ a list of $n$ claims $c_{1},\ldots,c_{n}$; batching claims reduces LLMq-calls (Appendix F.1)
-2:Partition $\{1,\ldots,n\}$ into disjoint subsets $S_{1},\ldots,S_{k}$, e.g. $S_{j}=\{i:i\mathop{\textrm{mod}}3=j-1\}$ for $j=1,2,3$
+2:Partition $\{c_{1},\ldots,c_{n}\}$ into disjoint subsets $S_{1},\ldots,S_{k}$, e.g. $S_{j}=\{c_{i}:i\mathop{\textrm{mod}}3=j-1\}$ for $j=1,2,3$
 3:for several rounds (e.g. 3), and in each round, for all $S_{j}$ in partition $(S_{1},\ldots,S_{k})$ do
@@
 5:     Provide LLMq with this modified list (but not $d$) and prompt it to recover the missing claims (Appendix F.2)
-6:     In the claims list $c_{1},\ldots,c_{n}$ replace the missing claims with the imputed claims
+6:     In the claims list $c_{1},\ldots,c_{n}$, replace the missing claims with the claims recovered in above step.
 7:end for
-8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only novel claims (Appendix F.9)
+8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only out-of-scope claims (Appendix F.11)
 9:Call LLMq to generate one short question per each novel claim, focusing on one key element of it (Appendix F.10)
@@
 3.1 Data Generation
-Here we describe our method of generating confusing questions for a given set of documents.
-We require each document to be short enough to fit in the LLM prompt along with instructions and examples, but long enough to make at least 5 to 10 separate claims (i.e.
+Here we describe our method of generating out-of-scope questions for a given set of documents.
+We require each document to be short enough to fit in the LLM prompt along with instructions and examples but long enough to make at least 4 to 10 separate claims (i.e.
 a few paragraphs long).
@@
 news articles published later than all pretraining cutoff dates, so that the LLMs cannot reproduce claims or answer questions correctly by pure hallucination.
-Thus, we collected 200 news articles published in 2024, each containing more than 150 words for each topic (Appendix E).
-Sentences were selected sequentially from the top until the total word count exceeded 300, using the NewscatcherBugara et al.
-(2020).
-Given a set of such documents, we generate multiple confusing questions about each document, along with some non-confusing ones for comparison.
-We start by picking LLMq (GPT-4o-mini by default in this paper), the “teacher” LLM that will generate the confusing questions.
-We prefer, but do not require, LLMq to be larger and stronger than the LLMr under evaluation.
+Thus, by using the Newscatcher Bugara et al.
+(2020), we collected 200 news articles published in 2024, each containing more than 150 words, for each topic (Appendix E).
+Sentences were selected sequentially from the top until the word count exceeded 300.
+Given a set of such documents, we generate multiple out-of-scope questions about each document and some in-scope ones for comparison.
+We start by picking LLMq (GPT-4o-mini by default in this paper), the “teacher” LLM that will generate the out-of-scope questions.
+We prefer but do not require LLMq to be larger and stronger than the LLMr under evaluation.
 For each document $d$, we call LLMq to generate a list of factual claims, then iteratively replace subsets of claims with hallucinated claims while maintaining semantic similarity to $d$ using guided hallucination, see Algorithm 1.
-We also generate some non-confusing questions for each document $d$ by calling LLMq either directly.
-3.3 Evaluation and Metrics
-We first assessed the quality of RAG-ConfusionQA by sampling a subset of questions and manually verifying whether they were indeed confusing or not (Section 3.4).
-Next, we evaluated how well various LLMr models could defuse the confusing questions in RAG-ConfusionQA.
-Given the large number of confusing questions, manual verification of LLMr responses is not feasible.
+We also generate in-scope questions for each document $d$ by calling LLMq directly.
+3.2 Evaluation
+We first assessed the quality of ScopeQA by sampling a subset of questions and manually verifying whether they were indeed out-of-scope or not (Section 3.3).
+Next, we evaluated how well various LLMr models could defuse the out-of-scope questions in ScopeQA.
+Given the large number of out-of-scope questions, manual verification of LLMr responses is not feasible.
 Therefore, we proposed AutoDefuseEval using GPT-4o-mini to automatically detect whether a response successfully defuses the question.
@@
 Question generation:
-What is the quality of confusing and non-confusing questions generated by Algorithm 1? (Section 3.4)
+What is the quality of out-of-scope and in-scope questions generated by Algorithm 1? (Section 3.3)
 Defusion detection:
-How accurately can AutoDefuseEval detect if an LLMr response to a confusing question defuses the confusion? (Section 3.4)
-Confusion response:
-How frequently does an LLMr, when responding to a confusing question using the standard prompt, generate a response that successfully defuses the confusion? (Section 4.1)
-Confusion detection:
-How accurately can LLMs identify which context-grounded questions are confusing and require special handling? (Section 4.2)
-We evaluate “question generation” and “defuse detection” through accuracy and confusion matrix.
-We only apply accuracy for “confusion response” and “confusion detection”.
-Algorithm A has the steps we run for each LLMr and for each $(q,d)$ pair.
-We run steps 1, 2, and 3 for “confusion detection” and 1, 4, 5, and 6 for “confusion response”.
+How accurately can AutoDefuseEval detect if an LLMr response to a out-of-scope question defuses the confusion? (Section 3.3)
+Out-of-scope response:
+How frequently does an LLMr, when responding to an out-of-scope question using different prompting methods, generate a response that successfully defuses the confusion? (Section 4.1)
+Out-of-scope detection:
+How accurately can LLMs identify which context-grounded questions are out-of-scope and require special handling? (Section 4.2)
+Appendix A has the steps we run for evaluating the above tasks.
+We run steps 1, 2, and 3 for “out-of-scope detection” and 1, 4, 5, 6, and 7 for “out-of-scope response”.
 Following the self-consistency method Wang et al.
 (2023a), we perform multiple LLM calls in steps  2 and  5 of Algorithm A, taking the majority vote to determine the final label.
-We also present the LLMq with few-shot exemplars of questions, responses, and explanations (Section 4.3).
-Model business entm food music news politics science sport tech travel Avg Std Dev
-GPT-3.5 18.90 12.39 13.55 15.48 15.15 17.13 10.81 $20.24^{\star}$ 17.12 10.01 15.08 3.21
-Llama 3.2 3B 68.34 66.11 64.89 67.90 67.50 66.94 57.98 $76.37^{\star}$ 67.07 60.06 66.31 4.70
-Mistral 7B v0.3 71.27 70.37 65.81 71.60 71.62 69.26 60.35 $79.00^{\star}$ 69.97 64.29 69.35 4.77
-Llama 3.1 70B 74.48 72.55 70.33 75.31 71.33 71.39 65.60 $\textbf{83.33}^{\star}$ 73.17 70.38 72.79 4.34
-Llama 3.1 8B 76.18 74.03 70.23 76.07 73.15 72.87 64.37 $81.45^{\star}$ 73.47 68.11 72.99 4.45
-Table 1: Evaluation of LLMs’ accuracy in defusing confusing questions from the RAG-ConfuseQA-Silver dataset across diverse news topics.
-For each LLM, the underscored value and stared value are the minimum and maximum values, respectively, across all the topics.
-Bold values are the maximum values across all the LLMs.
+Prompt Model business entm food music news politics science sport tech travel Avg Std Dev
+Basic GPT-3.5 18.06 10.96 12.32 14.27 13.59 15.95 10.15 $18.37^{\star}$ 15.59 9.63 13.89 2.97
+Llama 3.2 3B 63.59 61.55 61.09 62.07 63.49 64.01 53.35 $71.71^{\star}$ 63.69 56.51 62.11 4.60
+Mistral 7B v0.3 68.40 67.63 60.88 67.74 67.96 66.86 56.07 $76.23^{\star}$ 66.87 61.06 65.97 5.20
+Llama 3.1 8B 71.25 69.42 67.56 72.04 67.16 71.89 59.73 $77.90^{\star}$ 70.15 63.49 69.06 4.74
+Llama 3.1 70B 71.54 70.22 67.56 71.07 67.66 70.75 61.51 $79.76^{\star}$ 70.67 68.04 69.88 4.33
+Llama 3.3 70B 71.25 72.71 68.17 73.51 69.84 69.90 60.04 $\textbf{82.42}^{\star}$ 73.23 68.78 70.98 5.29
+Two-shot GPT-3.5 65.65 65.24 57.60 65.88 62.20 66.48 52.41 $71.71^{\star}$ 59.90 56.83 62.39 5.43
+Llama 3.2 3B 80.77 79.28 76.08 79.47 76.79 78.25 72.28 $84.09^{\star}$ 79.90 71.64 77.85 3.61
+Mistral 7B v0.3 80.86 77.69 74.54 78.79 79.17 79.11 67.47 $86.84^{\star}$ 78.15 71.43 77.40 5.02
+Llama 3.1 8B 83.32 80.98 80.18 84.16 78.77 82.53 72.70 $87.92^{\star}$ 82.77 78.41 81.17 3.87
+Llama 3.1 70B 85.18 83.37 80.60 85.83 79.76 82.24 75.84 $\textbf{90.57}^{\star}$ 83.08 79.79 82.62 3.84
+Llama 3.3 70B 85.48 81.27 80.39 84.26 79.27 82.34 75.84 $88.90^{\star}$ 83.49 78.62 81.99 3.57
+Zero-shot-CoT GPT-3.5 63.00 61.95 55.95 64.03 62.40 62.58 53.03 $71.71^{\star}$ 59.69 56.30 61.07 4.96
+Llama 3.2 3B 83.91 78.09 77.21 79.96 78.77 81.77 73.85 $85.46^{\star}$ 80.41 75.77 79.52 3.39
+Mistral 7B v0.3 77.72 78.39 70.94 78.59 79.07 77.02 67.26 $84.77^{\star}$ 76.41 72.06 76.22 4.69
+Llama 3.1 8B 79.39 78.09 74.85 78.40 78.67 79.20 70.29 $83.89^{\star}$ 77.64 74.50 77.49 3.43
+Llama 3.1 70B 90.28 87.85 87.27 88.56 86.51 89.08 83.68 $91.55^{\star}$ 86.97 85.19 87.69 2.21
+Llama 3.3 70B 93.52 91.24 90.55 92.57 90.77 92.12 89.12 $\textbf{95.09}^{\star}$ 93.85 90.90 91.97 1.71
+Table 2: Evaluation of LLMs’ accuracy in defusing out-of-scope questions from the ScopeQA-Silver dataset across diverse news topics.
+For each LLM, the underscored value and stared () value are the minimum and maximum values, respectively, across all the topics, within the same prompting method.
+Bold values are the maximum values for each topic across all the LLMs within the same prompting method.
 3.3 Human Annotation
-To evaluate “question generation” and “defuse detection” (Section 3.3), we recruited six computer science graduate students for manual labeling (Appendix B).
-As reflected in “Ground Truth Acc,” of Appendix B, Table  3, our method achieved 94.91% accuracy in generating confusing and non-confusing questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion.
+To evaluate “question generation” and “defusion detection” (Section 3.2), we recruited six computer science graduate students for manual labeling (Appendix B).
+As reflected in “Ground Truth Acc,” of Appendix B, Table  4, our method achieved 94.91% accuracy in generating out-of-scope and in-scope questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion.
 As shown in Appendix B, Figure 2, the confusion matrix on the left strongly agrees with the ground truth, with minimal errors (5 false negatives and 6 false positives).
@@
 4 Experimental Results
-4.1 Confusion Response
-We selected several open-source LLMs (Appendix 4) to evaluate their ability to generate responses that effectively defuse confusing questions with a standard prompt (Appendix F.6).
-As shown in Table 1, a) GPT-3.5 performs the worst among all six LLMs but with the smallest Std Dev across all topics, indicating consistent performance.
-On the other hand, Llama 3.1 70B and Llama 3.1 8B achieve comparable top performances with similar Std Dev, despite their differences in model size; b) most LLMs perform the worst in the “science” topic, except for GPT-3.5, for which “science” is the second worst.
-This is expected, as science topics often require more specialized domain knowledge to defuse confusion effectively; c) all LLMs achieve their best performance on the “sport” topic, demonstrating a clear trend.
-The strong performance across LLMs on the “sport” topic might be attributed to the relative simplicity and unambiguous nature of sports-related content.
-4.2 Confusion Detection
-By default, we use GPT-4o-mini as our LLMq for the task of “confusion detection,” which identifies whether the generated questions are confusing.
-In this section, we compare GPT-4o-mini with other LLMs.
-As illustrated in Figure 1, GPT-4o-mini achieves the best performance, while the smallest model, Llama 3.2 3B, performs the worst.
-Llama 3.2 3BMistral 7BLlama 3.1 8Bgpt-3.5Llama 3.1 70Bgpt-4o-mini606060608080808010010010010079.6379.6379.6379.6385.1985.1985.1985.1986.5786.5786.5786.5787.9687.9687.9687.9693.5293.5293.5293.5294.9194.9194.9194.91Different LLMsAccuracy (%)
-Figure 1: Evaluation of instruction-finetuned LLMs’ accuracy in classifying questions as confusing or not on the RAG-ConfusionQA-Gold dataset
+4.1 Out-of-Scope Response
+We evaluated open-source LLMs (Appendix 5) on their ability to defuse out-of-scope questions with three prompting methods (Appendix F.6,  F.7,  F.8).
+As shown in Table 2, a) Two-shot prompt boosts all evaluated LLMs’ accuracy substantially, especially GPT-3.5, likely due to the examples clarifying how to respond to out-of-scope questions.
+In contrast, Zero-shot-CoT Kojima et al.
+(2022) merely invites the LLM to “reason step by step” beats Two-shot on three out of six LLMs.
+b) Larger models (70B) benefit more from Zero-shot-CoT than from Two-shot prompting, achieving accuracy gains of 5.07% to 9.98%, while smaller models (3B to 8B) experience only marginal changes, ranging from -3.68% to 1.67%.
+This difference likely stems from larger models’ ability to better utilize their advanced reasoning capabilities and extensive knowledge; c) Most LLMs perform worst on the “science” topic and best on “sport”.
+This is expected, as scientific content tends to contain more implicit premises assumed by the author rather than explicitly stated facts, thus requiring more domain knowledge.
+In contrast, sports content is simpler and unambiguous, and the facts in the document are often clearly stated.
+4.2 Out-of-Scope Detection
+By default, we use GPT-4o-mini as our LLMq for “out-of-scope detection” task.
+In this section, we study the impact of different LLMs and methods.
+As illustrated in Figure 1, “Direct Generate” prompts the LLM to do the detection (Appendix F.3); “Unused-token Classification” represents a classifier trained on ScopeQA-Silver using the embedding of an appended unused-token (Appendix H) as a feature.
+Finally, “SBERT NLI” represents a classifier trained on ScopeQA-Silver using embeddings from a sentence-transformer model (Appendix H.3).
+Among these methods, GPT-4o-mini performs best with the “Direct Generate” approach.
+Notably, ScopeQA proves its utility by enabling even smaller models, like the Llama 3.2 3B-based classifier, to achieve accuracy comparable to much larger models, such as Llama 3.1 70B and GPT-4o-mini.
+To minimize interference with the LLM’s generation process, we append an unused-token, leaving its potential impact for future investigation.
+Figure 1: Evaluation of out-of-scope detection on ScopeQA-Gold.
 Module Accuracy
@@
 + examples 97.35
-+ self-consistency (n=3) 98.23
-+ self-consistency (n=9) 98.23
-Table 2: Evaluation of different prompts in the task “defusion detection” on RAG-ConfusionQA-Golden.
++ self-consistency (m=3) 98.23
++ self-consistency (m=9) 98.23
+Table 3: Evaluation of different prompts for “defusion detection” on ScopeQA-Gold.
 4.3 Ablation Study
-We examined the impact of various prompts on the task of “defusion detection.” As seen in Table 2, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $n=3$ further boosted performance to 98.23%.
-However, increasing the number of generated sequences beyond three did not lead to additional improvements.
-Since RAG-ConfusionQA-Golden is a smaller dataset, using $n=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger RAG-ConfusionQA-Silver dataset.
-Therefore, we adopted $n=9$ in our experiments to balance accuracy and expenses (Appendix F.10).
+We examined the impact of various prompts on “defusion detection” task.
+As seen in Table 3, incorporating examples increased accuracy from 87.61% to 97.35%, and applying self-consistency with $m=3$ further boosted performance to 98.23%.
+However, further increasing $m$ did not lead to additional improvements.
+Since ScopeQA-Gold is a smaller dataset, using $m=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger ScopeQA-Silver dataset.
+Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix G).
 5 Conclusion
-We introduced RAG-ConfusionQA, a benchmark designed to assess how effectively an LLM defuses confusion when responding to confusing questions using a standard prompt.
-Through human annotation, we demonstrated that RAG-ConfusionQA achieves a 94.91% accuracy in generating confusing and non-confusing questions, while our proposed AutoDefusionEval achieves a 98.23% accuracy in detecting whether an LLM’s response defuses confusion.
-Additionally, our evaluation highlights that current LLMs often attempt to answer confusing questions to varying degrees.
-For future work, we aim to develop methods to better detect confusing questions and address the tendency of LLMs to respond to them directly instead of defusing the confusion.
+We introduced ScopeQA, a benchmark for evaluating LLMs’ ability to defuse out-of-scope questions in RAG.
+Human annotation validates that ScopeQA achieves 94.91% accuracy in distinguishing out-of-scope questions, while our AutoDefusionEval achieves 98.23% accuracy in detecting effective defusion responses.
+We further demonstrated ScopeQA’s utility by training out-of-scope detectors.
+For future work, we plan to investigate the impact of appending an unused-token on the generation process.
+Additionally, we aim to explore instruction fine-tuning by training LLMs on ScopeQA and evaluating their performance.
+We also intend to apply for OpenAI funding to leverage more powerful models, such as GPT-4o, to enhance data quality.
 6 Acknowledgement
@@
 7 Limitations
-In this paper, we only evaluated LLMs’ abilities to generate responses that effectively defuse confusing questions when prompted with a standard prompt (Appendix F.6).
-We chose this prompt because, ideally, LLMs should be able to address confusion with a basic, straightforward prompt.
-However, we have not yet assessed the performance of LLMs across different prompt variations.
+In this paper, we evaluated LLMs’ abilities to defuse out-of-scope questions using three prompting methods: basic prompt, Two-shot prompt, and Zero-shot-CoT prompt.
+These methods were chosen because LLMs should ideally handle out-of-scope questions with straightforward prompts.
+However, we did not assess the performance of LLMs fine-tuned on ScopeQA-Silver due to computational resource constraints.
+Additionally, we empirically appended an unused-token to the end of the prompt for classification while minimizing its influence on LLMs’ generation processes.
+The effect of appending such tokens on generation remains unexplored and is left for future work.
+Finally, budget constraints prevented us from using more powerful models, such as GPT-4o or GPT-o1, which could theoretically produce higher-quality datasets.
+Future users can adopt our framework with more advanced models and alternative data sources to generate even higher-quality data tailored to their needs.
 Appendix A Algorithm
 Algorithm 2 Evaluation steps
-1:Call LLMr on $(d,q)$ with the standard RAG prompt and get its response $r$ (Appendix F.6)
-2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is confusing given $d$ (Appendix F.3)
+1:Call LLMr on $(d,q)$ with RAG prompt and get its response $r$ (Appendix F.6,  F.7,  F.8)
+2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is out-of-scope given $d$ (Appendix F.3)
 3:Aggregate $m$ predictions by majority vote
-4:if $q$ is generated as confusing then
+4:if $q$ is generated as out-of-scope then
 5:     Call LLMq $m$ times on $(d,q,r)$ to check and explain if $r$ defused the confusion (Appendix F.4)
@@
 Appendix B Human Annotation
-Figure 2: Confusion matrix of confusion and defusion on RAG-ConfusionQA-Golden.
+Figure 2: Confusion matrix of out-of-scope and defusion on ScopeQA-Gold.
 Group Name Cohen’s Kappa Annotator Acc Group’s Agree Acc Ground Truth Acc
-confusion defusion confusion defusion confusion defusion confusion defusion
+Out-of-Scope defusion Out-of-Scope Defusion Out-of-Scope Defusion Out-of-Scope Defusion
 1   — 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
@@
   — 98.61 100.00
-Table 3: Agreement scores and the performance of our method on the RAG-ConfusionQA-Golden.
-Following the methodology from Cochran (1977), representing approximately 20,000 synthetic questions in RAG-ConfuseQA at a 95% confidence level, with a 4% margin of error, and an anticipated accuracy of $90\pm 5\%$, required labeling at least 202 questions.
-We sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators.
-As shown in Table 3, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is confusing and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency.
+Table 4: Agreement scores and the performance of our method on the ScopeQA-Gold.
+Following the methodology from Cochran (1977), we sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators.
+As shown in Table 4, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is out-of-scope and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency.
 Two more annotators manually resolved all disagreements to establish ground truth labels.
-Six annotators are computer science graduate students recruited via email and incentivized with complimentary meals.
-Another two annotators are the authors of this paper, who are also majoring in computer science.
+All the annotators are computer science graduate students recruited via email and incentivized with complimentary meals.
+All annotators were trained by first reviewing the annotation guidelines and then labeling a small dataset previously labeled by the authors.
+The annotation guideline and annotation tool will be released along with the data and code upon acceptance.
 We used two metrics to assess our method: “Annotator Acc” and “Group’s Agree Acc.” “Annotator Acc” treats each annotator’s labels as ground truth, computing the accuracy of our method’s generated labels.
@@
 Llama 3.1 70B Instruct Turbo FP8 2023/12
+Llama 3.3 70B Instruct Turbo FP8 2023/12
 Mistral (7B) Instruct v0.3 FP16 2024/5
 gpt-3.5-turbo   — 2021/9
-Table 4: Different LLMr to be evaluated. In RAG-ConfuseQA-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cut-off date 5/22/2024.
-Appendix D Confusing Types
+Table 5: Different LLMr to be evaluated. In ScopeQA-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cut-off date 5/22/2024.
+Appendix D Out-of-scope Example
 Document:
@@
 Jefferson is an automatic start whenever he’s healthy, and his star power may be needed more on certain fantasy rosters in Week 3 depending on how many injuries affect a single roster.
-Question 1: How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?
-This question is confusing as it asks about "Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft" which was never discussed in the document. So this belongs to Not Mention
-Question 2: How will Justin Jefferson’s struggles with hamstring issues in Week 3 affect the Vikings’ chances of winning against the Houston Texans?
-This question assumes that Justin Jefferson is dealing with hamstring issues, which is not mentioned in the document.
-The document only refers to a quad injury that Jefferson had, which is no longer a concern as of Week 3.
-The document makes it clear that Jefferson is healthy and will play without any restrictions.
-Therefore, the question is based on a False Premise
-Question 3: How will his performance impact the team’s chances this season, considering they are missing key players?
-This question uses the Ambiguous pronoun “his” without specifying who it refers to, making it unclear whether the question is asking about Justin Jefferson, CeeDee Lamb, or another player mentioned in the document.
-Additionally, “the team” is not defined explicitly, as both the Minnesota Vikings and other teams are referenced in the document.
-Without clarifying which player or team the question is referring to, it is impossible to provide an accurate answer based on the document.
-Appendix E Statistics of RAG-ConfusionQA
-Attr Golden Silver
-# of documents 45 2,000
-# of non-confusing questions 103 9,052
-# of confusing questions 113 10,218
+Question: How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?
+Explanation: This question is out-of-scope as it asks about "Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft" which was never discussed in the document.
+Appendix E Statistics of ScopeQA
+Attr Gold Silver
+# of documents 45 1955
+# of in-scope questions 103 8,949
+# of out-of-scope questions 113 10,105
 average words per document 280 276
-average words per non-confusing question 16 16
-average words per confusing question 16 16
-Table 5: Statistics of RAG-ConfusionQA. Golden is sampled from Silver and annotated by annotators (Appendix B).
+average words per in-scope question 16 16
+average words per out-of-scope question 16 16
+Table 6: Statistics of ScopeQA. Gold is sampled from crawled 2000 news and annotated by annotators (Appendix B).
 Appendix F Prompt
 F.1 Extract Claims Prompt
-{mdframed}
-You will be provided with a document delimited by triple quotes.
-Read the document and follow user’s instructions.
-Read the document and list {num_fact} most important facts it contains. Each fact should be
-stated in a clear, standalone sentence with sufficient context to be understood
-independently, avoiding undefined pronouns. Ensure that each fact is directly derived from
-the document and does not include any information not mentioned within it.
+You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
+Read the document and list {num_fact} most important facts it contains.
+Each fact should be stated in a clear, standalone sentence with sufficient context to be understood independently, avoiding undefined pronouns.
+Ensure that each fact is directly derived from the document and does not include any information not mentioned within it.
 Document:
 """{document}"""
-{num_fact} most important facts:"
+{num_fact} most important facts:
 F.2 Recover Missing Claims Prompt
-{mdframed}
-Read the document below with a list of {num_fact} facts it contains. Note that some of the
-facts are missing represented by (missing). Your task is to guess the missing facts could
-have said and complete the missing facts. Each fact should be stated in a clear, standalone
-sentence with sufficient context to be understood independently, avoiding undefined pronouns.
-Please complete the missing facts and return all the {num_fact} facts in original order. You
-must return all the {num_fact} facts.
+Read the document below with a list of {num_fact} facts it contains.
+Note that some of the facts are missing represented by (missing).
+Your task is to guess the missing facts could have said and complete the missing facts.
+Each fact should be stated in a clear, standalone sentence with sufficient context to be understood independently, avoiding undefined pronouns.
+Please complete the missing facts and return all the {num_fact} facts in original order.
+You must return all the {num_fact} facts.
 Document:
 """document"""
-Completed list of facts:"
-F.3 Confusion Judgement Prompt
-{mdframed}
-You will be provided with a document delimited by triple quotes.
-Read the document and follow user’s instructions.
-Read the document and the question. Check if the question contains any confusing
-parts, false assumptions, incorrect information, or mentions of people, locations,
-organizations, or other entities that are not present in the document.
+Completed list of facts:
+F.3 Out-of-scope Judgement Prompt
+You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
+Read the document and the question. Check if the question contains any confusing parts: mentions of people, locations, organizations, or other entities that are not present in the document
 Document:
@@
 {question}
-Does the question contain any confusing parts, false assumptions, incorrect
-information, or mentions of people, locations, organizations, or other entities
-that are not present in the document? Please start by explaining your reasoning.
+Does the question contain any confusing parts: mentions of people, locations, organizations, or other entities that are not present in the document?
+Please start by explaining your reasoning.
 End your response with ’The answer is: Yes or No.’:
-F.4 Defusion Judgement Prompt
-{mdframed}
-You will be provided with a document delimited by triple quotes.
-Read the document and follow user’s instructions.
-Read the document and answer the question based on the document.
+F.4 Response Defusion Judgement Prompt
+You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions. Read the document and answer the question based on the document.
 Document:
@@
 Question:
-    {example_question}
-{defuse_answer_examples}
-{Defusion Instruction}
-{defuse_reasoning_examples}
+{example_questions}
+Answer:
+{example_LLM_responses}
+{Defusion Instruction (Appendix  F.5)}
+{example_defuse_reasoning}
 Read the document and answer the question based on the document.
@@
 {question}
-{response_from_gpt3.5}
-{Defusion Instruction}
+Answer:
+{LLM_response}
+{Defusion Instruction (Appendix  F.5)}
 F.5 Defusion Instruction
-{mdframed}
 The question presented is confusing and unanswerable based on the document provided.
-Read the response carefully. If the response attempts to answer the question,
-like answering ’Yes’ to agree the question or ’No’ to disagree the question,
-you should immediately conclude with The answer is: No. even if it later
-acknowledges the questions confusion. Otherwise, determine if the response
-clarifies that the question cannot be answered or confusing or matches any of the
-issues listed:
-    False Premise: The response points out that the question is based
-    on an incorrect assumption.
-    Not Mentioned: The response indicates that the information needed to answer
-    the question is not available or not mentioned in the document.
-    Ambiguous: The response points out unclear pronouns or vague terms in the
-    question that prevent a clear answer.
-Begin with your reasoning and finally conclude with The answer is: Yes. or
-The answer is: No.. Note that if the response attempts to answer the question
-at any point, like answering ’Yes’ to agree the question or ’No’ to disagree the
-question, you should immediately conclude with The answer is: No. even if it
-later acknowledges the questions confusion:
-F.6 RAG Prompt
-{mdframed}
-You will be provided with a document delimited by triple quotes.
-Read the document and follow user’s instructions.
+Read the response carefully.
+If the response attempts to answer the question, like answering ’Yes’ to agree with the question or ’No’ to disagree with the question, you should immediately conclude with ‘The answer is: No.’ even if it later acknowledges the question’s confusion.
+Otherwise, determine if the response clarifies that the question cannot be answered or if the response indicates that the information needed to answer the question is not available or not mentioned in the document.
+Begin with your reasoning and finally conclude with ‘The answer is: Yes.’ or ‘The answer is: No.’.
+Note that if the response attempts to answer the question at any point, like answering ’Yes’ to agree with the question or ’No’ to disagree with the question, you should immediately conclude with ‘The answer is: No.’ even if it later acknowledges the question’s confusion:
+F.6 RAG Basic
+You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
 Read the document and answer the question based on the document.
@@
 {question}
-F.7 Non-confuing Question Generation Prompt
-{mdframed}
+Answer:
+F.7 RAG Two-shot
+You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
+Read the document and answer the question based on the document.
+Document:
+"""{example_document}"""
+Question:
+{two_example_questions}
+Answer:
+{two_example_responses}
+Read the document and answer the question based on the document.
+Document:
+"""{document}"""
+Question:
+{question}
+Answer:
+F.8 RAG Zero-shot-CoT
+You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
+Read the document and reason step by step to answer the question based on the document. If the question cannot be answered using the document, state explicitly that the question cannot be answered.
+Document:
+"""{document}"""
+Question:
+{question}
+Answer:
+F.9 In-scope Question Generation Prompt
 You will be provided with a document delimited by triple quotes.
 Read the document and follow user’s instructions.
-Read the document attentively and compile a numbered list of the top {num_q} questions
-that the document directly answers. Ensure each question is clear, accurate, and devoid
-of confusion, false assumptions, undefined pronouns, or misinformation. Avoid referencing
-people, locations, organizations, or other entities not explicitly mentioned in the
-document. Construct each question to be thought-provoking, containing between 13 to 18
-words, and sufficiently detailed to avoid being overly straightforward.
+Read the document attentively and compile a numbered list of the top {num_q} questions that the document directly answers.
+Ensure each question is clear, accurate, and devoid of confusion, false assumptions, undefined pronouns, or misinformation.
+Avoid referencing people, locations, organizations, or other entities not explicitly mentioned in the document.
+Construct each question to be thought-provoking, containing between 13 to 18 words, and sufficiently detailed to avoid being overly straightforward.
 Document:
@@
 Questions:
-F.8 Confusing Question Generatioin Prompt
-{mdframed}
-You will be provided with a document delimited by triple quotes.
-Read the document and follow user’s instructions.
-Read the document and review the list of hallucinated facts. For each hallucinated fact,
-craft a single, specific and concise question containing 13 to 18 words that incorporates
-the key element of the fact, ensuring the question is intentionally confusing.
+F.10 Out-of-scope Question Generation Prompt
+You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
+Read the document and review the list of hallucinated facts.
+For each hallucinated fact, craft a single, specific and concise question containing 13 to 18 words that incorporate the key element of the fact, ensuring the question is intentionally confusing.
 The question should not be answerable using any information present in the document.
-The question should not combine multiple queries and each question should address only
-one specific aspect. If a question cannot be formulated for a particular hallucinated
-fact, you may omit it.
+The question should not combine multiple queries and each question should address only one specific aspect.
+If a question cannot be formulated for a particular hallucinated fact, you may omit it.
 Document:
@@
 Questions:
-F.9 Remove Claims Prompt
-{mdframed}
-You will be provided with a document delimited by triple quotes.
-Read the document and follow user’s instructions.
-Read the document below with a list of {num_true_fact} ground-truth facts it contains
-and a list of {num_false_fact} hallucinated facts that are not supported by the document.
-Your task is to remove any hallucinated facts that can be supported by either the
-document or the {num_true_fact} ground-truth facts.
+F.11 Remove Claims Prompt
+You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
+Read the document below with a list of {num_true_fact} ground-truth facts it contains and a list of {num_false_fact} hallucinated facts that are not supported by the document.
+Your task is to remove any hallucinated facts that can be supported by either the document or the {num_true_fact} ground-truth facts.
 Please only return the remaining hallucinated facts, along with their original order numbers.
@@
 """{example_document}"""
-4 ground-truth facts:
+{num_true_fact} ground-truth facts:
 {true_facts}
@@
 {remained_facts}
-Read the document below with a list of {num_true_fact} ground-truth facts it contains
-and a list of {num_false_fact} hallucinated facts that are not supported by the document.
-Your task is to remove any hallucinated facts that can be supported by either the
-document or the {num_true_fact} ground-truth facts.
+Read the document below with a list of {num_true_fact} ground-truth facts it contains and a list of {num_false_fact} hallucinated facts that are not supported by the document.
+Your task is to remove any hallucinated facts that can be supported by either the document or the {num_true_fact} ground-truth facts.
 Please only return the remaining hallucinated facts, along with their original order numbers.
@@
 """{document}"""
-{num_true_fact} ground-truth facts:
+{num_true_fact} ground-truth facts
 {true_facts}
@@
 Remaining hallucinated facts:
-F.10 Expenses
-Generating the RAG-ConfusionQA dataset using GPT-4o-mini costs approximately \$50.
+Appendix G Expenses
+Generating the ScopeQA dataset using GPT-4o-mini costs approximately \$80.
 For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/.
-Excluding the dataset generation, all other experiments cost around \$60.
+Excluding the dataset generation, all other experiments cost around \$100.
+Appendix H Embedding Classifier
+H.1 Implementation Details
+We train the classifier on ScopeQA-Silver, which is split as: 80% for training, 10% for evaluation, and 10% for testing. After training, we test the classifier on ScopeQA-Gold data.
+We append the unused-token ID at the end of the tokenized input sequence.
+After the LLM processes the input, we extract the unused-token’s hidden state from the model’s last hidden layer output.
+Since this token is not part of the LLM’s training and does not carry predefined meanings, it serves as a placeholder to aggregate information from the model’s attention over all preceding tokens in the sequence.
+As the final token in the input, its embedding encapsulates the model’s representation of the entire input sequence, reflecting the information inferred from the input sequence.
+The unused-token’s hidden state is used as input features for a 2-layer MLP, which is optimized with a BCE loss.
+A dropout of 0.1 is applied after the first layer.
+We use Adam optimizer with 1e-4 learning rate.
+We train each classifier for 10 epochs with batch size 8.
+We save the best validation checkpoint and report its performance on ScopeQA-Gold.
+H.2 Unused-Token
+For the Llama family, we use the <|reserved_special_token_0|> as the unused-token.
+This token is not trained and is reserved for future use cases such that the users do not need to resize the vocabulary.
+For Mistral-7B-v0.3, we use [control_555] as the unused-token because Mistral’s official documentation indicated that they reserve 768 control tokens for future use, and none of them are used during training.
+555 is a random choice.
+H.3 Sentence BERT NLI baseline
+To create a baseline classification performance, we adapt NLI style training using an off-the-shelf sentence-transformers/all-mpnet-base-v2 model from HuggingFace.
+Specifically, we follow Reimers and Gurevych (2019) to extract document and question [CLS] embeddings $u$ and $v$ to use $[u;v;|u-v|]$ as the input feature for the 2-layer MLP to classify.

### v3

@@
-ScopeQA: A Framework for Generating Out-of-Scope Questions for RAG
-Abstract
-Conversational AI agents use Retrieval Augmented Generation (RAG) to provide verifiable document-grounded responses to user inquiries.
-However, many natural questions do not have good answers: about 25% contain false assumptions Yu et al.
-(2023), and over 50% are ambiguous Min et al.
-(2020).
-RAG agents need high-quality data to improve their responses to confusing questions.
-This paper presents a novel guided hallucination-based method to efficiently generate a diverse set of borderline out-of-scope confusing questions for a given document corpus.
-We conduct an empirical comparative evaluation of several large language models as RAG agents to measure the accuracy of confusion detection and appropriate response generation.
-We contribute a benchmark dataset to the public domain.
-ScopeQA: A Framework for Generating Out-of-Scope Questions for RAG
-Zhiyuan Pengthanks: Equal contribution.
-Santa Clara University Santa Clara, CA, USA zpeng@scu.edu                        Jinming Nian11footnotemark: 1 Santa Clara University Santa Clara, CA, USA jnian@scu.edu                        Alexandre Evfimievski aevfim@outlook.com                        Yi Fang Santa Clara University Santa Clara, CA, USA yfang@scu.edu
-1 Introduction
-Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents Lewis et al.
-(2020); Guu et al.
-(2020); Gao et al.
-(2024).
+ELOQ: Resources for Enhancing LLM Detection of Out-of-Scope Questions
+Abstract.
+Large Language Models (LLMs) are widely used in Conversational AI systems to generate responses to user inquiries.
+However, many natural questions lack well-defined answers.
+While existing studies primarily focus on question types such as false premises, they often overlook out-of-scope questions, where the provided document is semantically highly similar to the query but does not contain the required answer.
+In this paper, we propose a guided hallucination-based method to efficiently generate a diverse set of out-of-scope questions from a given document corpus.
+We then evaluate multiple LLMs based on their effectiveness in confusion detection and appropriate response generation.
+Furthermore, we introduce an improved method for detecting such out-of-scope questions, enhancing the reliability of LLM-based question-answering systems.
+Large Language Models, Question Answering, Out-of-Scope Question
+*Both authors contributed equally to this work.
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processing
+1. Introduction
+Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents (Lewis et al., 2020; Guu et al., 2020; Gao et al., 2024).
 Given an inquiry, a RAG system retrieves a relevant document from a curated knowledge base and provides it to a large language model (LLM) that generates a response.
-RAG performance is evaluated on dimensions like the response accuracy, faithfulness, completeness, answer and context relevance, and style Es et al.
-(2024); Saad-Falcon et al.
-(2024); Liu et al.
-(2024); Zheng et al.
-(2023).
-Errors may stem from a document-to-question mismatch, hallucinating parts of the response, wrong understanding of the document or the question.
-Many user questions have no direct answer: some may rely on a false premise, some may be ambiguous in what they ask, and some are out of scope for the knowledge base.
-A well-designed RAG system should detect such confusing questions and respond appropriately, either by clarifying the confusion or seeking help.
-Dataset RAG Setting Source Category Cut-off-knowledge Annotation Method
-AMBIGQA Min et al. (2020) Wiki Multi-answers Human
-CREPE Yu et al. (2023) Reddit False Premise Human
-FalseQA Hu et al. (2023)  False Premise Human
-KG-FPQ Zhu et al. (2024) Wiki False Premise LLM
-FAITH Yuan et al. (2024) Wiki False Premise LLM
-FLUB Li et al. (2024a) Baidu Tieba Cunning Texts Human
-ScopeQA News Out-of-scope Human+LLM
-Table 1: ScopeQA vs Existing Datasets
+RAG performance is evaluated on dimensions such as response accuracy, faithfulness, completeness, answer and context relevance, and style (Es et al., 2024; Saad-Falcon et al., 2024; Liu et al., 2024; Zheng et al., 2023).
+Errors may stem from a document-to-question mismatch, hallucinated parts of the response, or misunderstanding of the document or the question.
+Many user questions have no direct answer: some rely on a false premise, others are ambiguous in their intent, and some are out-of-scope for the knowledge base.
+Prior studies indicate that about 25% of natural questions contain false assumptions (Yu et al., 2023) and over 50% are ambiguous (Min et al., 2020a).
+Additionally, out-of-scope errors have been identified as the leading cause of enterprise AI assistants generating responses that appear convincing but are, in fact, incorrect (Maharaj et al., 2024).
+A well-designed RAG system should implement an exception-handling mechanism that detect confusing questions and responds by clarifying the confusion, retrieving additional documents, or seeking assistance.
+Dataset Source Category Cutoff knowledge Annotation Method
+AMBIGQA (Min et al., 2020a) Wiki Multi-answers Human
+CREPE (Yu et al., 2023) Reddit False Premise Human
+FalseQA (Hu et al., 2023)  False Premise Human
+KG-FPQ (Zhu et al., 2024) Wiki False Premise LLM
+FAITH (Yuan et al., 2024) Wiki False Premise LLM
+FLUB (Li et al., 2024a) Baidu Tieba Cunning Texts Human
+ELOQ News Out-of-scope Human+LLM
+Table 1. ELOQ vs Existing Datasets
 Enabling “deconfusion” in a RAG system involves three challenges: (1) creating a diverse dataset of confusing questions, (2) training a classifier to detect confusion, and (3) developing a response generator for each confusion type.
-As shown in Table 1, much prior work focuses on false-premise questions that contradict pretraining knowledge but neglects out-of-scope queries, which lack answers in retrieved documents.
-Borderline out-of-scope questions can confuse LLMs into making up a hallucinated answer.
-This type of confusion is the focus of this paper.
-While queries in MS MARCO Nguyen et al.
-(2016) may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cut-off.
-ScopeQA fills this gap by providing a diverse benchmark for out-of-scope questions, leveraging LLM-assisted generation to minimize human annotation effort.
+As shown in Table 1, much prior work focuses on false-premise questions that contradict pretraining knowledge, but often neglects out-of-scope queries for which no answers can be found in retrieved documents.
+Out-of-scope questions can mislead LLMs into generating hallucinated answers that appear correct, but are incorrect, which represents the most dangerous type of error (Maharaj et al., 2024).
+Our focus is to mitigate this type of confusion.
+This paper proposes a framework for automatically generating out-of-scope questions and evaluates its effectiveness in training classifiers to detect out-of-scope situations.
+While queries in MS MARCO (Nguyen et al., 2016) may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cutoff.
+ELOQ fills this gap by providing a diverse benchmark for out-of-scope questions, leveraging LLM-assisted generation to reduce human annotation effort.
 Our contributions are as follows:
-  • We propose and implement a method to generate borderline out-of-scope questions for given documents efficiently.
-  • We validate our synthetic data by training out-of-scope question detectors.
-  • We evaluate LLMs as RAG agents for detecting confusion and generating responses.
-  • Our code and dataset are released at https://github.com/zhiyuanpeng/ScopeQA.git.
-2 Related Work
-Benchmark Datasets
-Most of the recent work on LLM responses to confusing questions focuses on stand-alone questions answered using LLM’s parametric knowledge without a context document.
-The questions may be naturally collected from online sources Li et al.
-(2024b); Yu et al.
-(2023), written by human annotators Hu et al.
-(2023), or synthetically generated Zhu et al.
-(2024); Yuan et al.
-(2024).
-A comparison of these released benchmark datasets is shown in Table 1.
-Question Generation
-Zhu et al.
-(2024); Yuan et al.
-(2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of subject + relation + object from Wikipedia, replacing the object with a similar but incorrect one, and filling in a template.
-Content-grounded (non-confusing) synthetic data generation is studied in Yehudai et al.
-(2024); Wang et al.
-(2023b); Lee et al.
-(2023).
-Given the grounding passages, they generate data entries by few-shot prompting with task-specific examples, then filter to ensure data quality, faithfulness, and diversity.
-Mitigation
-is commonly achieved with in-context learning  Zhu et al.
-(2024); Li et al.
-(2024b), chain-of-thought (CoT)  Wei et al.
-(2022) and fine-tuning.
-Experiments show that fine-tuning outperforms the other two methods Hu et al.
-(2023).
-CoT yields inconsistent improvements, while in-context learning improves performance with more examples Li et al.
-(2024b).
+  • We propose and implement an efficient method for generating synthetic out-of-scope questions from a given set of documents.
+  • We demonstrate the utility of our synthetic data in training out-of-scope exception detectors.
+  • We conduct a comparative evaluation of multiple LLMs as RAG agents to measure their confusion detection accuracy.
+  • Our code  [^1] https://github.com/zhiyuanpeng/ELOQ.git and dataset  [^2] https://huggingface.co/datasets/zhiyuanpeng/ELOQ are made publicly available.
+2. Related Work
+2.1. Benchmark Datasets
+Recent work on LLM responses to confusing questions primarily focuses on stand-alone questions answered using the LLM’s pretraining knowledge, without a supporting context document.
+These questions may be naturally collected from online sources (Li et al., 2024b; Yu et al., 2023), written by human annotators (Hu et al., 2023), or synthetically generated (Zhu et al., 2024; Yuan et al., 2024).
+Released benchmark datasets with confusing questions include KG-FPQ (Zhu et al., 2024), FalseQA (Hu et al., 2023), FLUB (Li et al., 2024b), and CREPE (Yu et al., 2023).
+Earlier research on ambiguous open-domain questions (Min et al., 2020b) introduced AmbigNQ, a dataset of naturally ambiguous questions.
+Table 1 presents a comparison of these benchmark datasets.
+2.2. Question Generation
+(Zhu et al., 2024; Yuan et al., 2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of (subject, relation, object) using Wikipedia data, replacing the object with a similar but incorrect one, and then filling in a template to generate a false premise question.
+Content-grounded (non-confusing) synthetic data generation has been studied in (Yehudai et al., 2024; Wang et al., 2023a; Lee et al., 2023).
+Given the grounding passages, these methods generate data entries using few-shot prompting with task-specific examples, followed by filtering to ensure data quality, faithfulness, and diversity.
+2.3. Mitigation
+Mitigation is commonly achieved through in-context learning  (Zhu et al., 2024; Li et al., 2024b), chain-of-thought (CoT)  (Wei et al., 2022) and fine-tuning.
+Experiments show that fine-tuning outperforms both in-context learning and CoT reasoning (Hu et al., 2023).
+CoT reasoning leads to inconsistent improvements, whereas in-context learning improves performance as the number of examples increase (Li et al., 2024b).
 Yuan et al.
-(2024) identifies  1% of attention heads responsible for confusion by false-premise questions, showing a 20% performance gain by constraining these heads.
+(2024) identifies approximately 1% of attention heads as responsible for confusion caused by false-premise questions and demonstrates a 20% performance gain by constraining these heads.
 Kim et al.
 (2021) retrieves a Wikipedia page for each question and responds with the identified unverifiable premises if the page does not entail them.
-3 Methods
-We developed the ScopeQA data generator to evaluate and mitigate LLM confusion when responding to out-of-scope questions.
-We assume that the retrieved document is the best available match for the question and focus on the response generation step of the RAG pipeline.
-In this step, given a standard prompt $p$, an out-of-scope question $q$, and a retrieved document $d$, the LLMr generates a response $r$.
-Ideally, $r$ should clarify the confusing part of the question, a process we refer to as defusion (or de-confusion), rather than attempting to answer the question directly and hallucinate.
 Algorithm 1 The guided hallucination method
-1:Call LLMq to extract from $d$ a list of $n$ claims $c_{1},\ldots,c_{n}$; batching claims reduces LLMq-calls (Appendix F.1)
+1:Call LLMq to extract from $d$ a list of $n$ claims $c_{1},\ldots,c_{n}$; batching claims reduces LLMq-calls (Appendix B.1)
 2:Partition $\{c_{1},\ldots,c_{n}\}$ into disjoint subsets $S_{1},\ldots,S_{k}$, e.g. $S_{j}=\{c_{i}:i\mathop{\textrm{mod}}3=j-1\}$ for $j=1,2,3$
@@
 4:    In the claims list $c_{1},\ldots,c_{n}$ replace all claims $c_{i}$ where $i\in S_{j}$ by text “(missing)
-5:     Provide LLMq with this modified list (but not $d$) and prompt it to recover the missing claims (Appendix F.2)
-6:     In the claims list $c_{1},\ldots,c_{n}$, replace the missing claims with the claims recovered in above step.
+5:    Provide LLMq with this modified list (but not $d$) and prompt it to recover the missing claims (Appendix B.2)
+6:    In the claims list $c_{1},\ldots,c_{n}$, replace the missing claims with the claims recovered in the above step.
 7:end for
-8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only out-of-scope claims (Appendix F.11)
-9:Call LLMq to generate one short question per each novel claim, focusing on one key element of it (Appendix F.10)
-10:Call LLMq to filter out any questions answerable in the context of $d$ (Appendix F.3)
-3.1 Data Generation
-Here we describe our method of generating out-of-scope questions for a given set of documents.
+8:Call LLMq to remove all claims supported by $d$ or by the original claims, leaving only out-of-scope claims (Appendix B.11)
+9:Call LLMq to generate one short question per each novel claim, focusing on one key element of it (Appendix B.10)
+10:Call LLMq to filter out any questions answerable in the context of $d$ (Appendix B.3)
+3. Methods
+We developed the ELOQ data generator to evaluate and mitigate LLM confusion when responding to out-of-scope questions for a given document.
+We assume that the user has limited knowledge of the document, resulting in asking questions that appear relevant but are actually out-of-scope and cannot be answered based on the document’s content.
+In this step, given a standard prompt $p$, an out-of-scope question $q$, and a relevant document $d$, the LLMr generates a response $r$.
+Ideally, $r$ should clarify the confusing part of the question – a process we refer to as defusion (or de-confusion) – rather than attempting to answer the question directly and risking hallucination.
+Algorithm 2 Evaluation steps
+1:Call LLMr on $(d,q)$ with RAG prompt and get its response $r$ (Appendix B.6,  B.7,  B.8)
+2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is out-of-scope given $d$ (Appendix B.3)
+3:Aggregate $m$ predictions by majority vote
+4:if $q$ is generated as out-of-scope then
+5:     Call LLMq $m$ times on $(d,q,r)$ to check and explain if $r$ defused the confusion (Appendix B.4)
+6:     Aggregate $m$ predictions by majority vote
+7:end if
+Attr Gold Silver
+# of documents 45 1955
+# of in-scope questions 103 8,949
+# of out-of-scope questions 113 10,105
+average words per document 280 276
+average words per in-scope question 16 16
+average words per out-of-scope question 16 16
+Table 2. Statistics of ELOQ. Gold is sampled from crawled 2000 news and annotated by annotators (Section 3.4).
+3.1. Data Collection
+We focus on the scenario where the domain knowledge is maintained separately from the LLM in a document database.
+Hence, we prefer documents that are novel to the evaluated LLMs, such as news articles published after all pretraining cutoff dates, ensuring that LLMs cannot reproduce claims or answer questions correctly by pure hallucination.
+Instead, they must generate responses grounded in their understanding of the news content.
+Therefore, using the Newscatcher (Bugara et al., 2020), we collected 200 news articles published between Jan 1, 2024, and September 30, 2024, for each topic333sport, business, science, food, politics, travel, entertainment, music, news, tech.
 We require each document to be short enough to fit in the LLM prompt along with instructions and examples but long enough to make at least 4 to 10 separate claims (i.e.
 a few paragraphs long).
-We prefer documents that are novel to the LLMs under evaluation, e.g.
-news articles published later than all pretraining cutoff dates, so that the LLMs cannot reproduce claims or answer questions correctly by pure hallucination.
-Thus, by using the Newscatcher Bugara et al.
-(2020), we collected 200 news articles published in 2024, each containing more than 150 words, for each topic (Appendix E).
-Sentences were selected sequentially from the top until the word count exceeded 300.
-Given a set of such documents, we generate multiple out-of-scope questions about each document and some in-scope ones for comparison.
-We start by picking LLMq (GPT-4o-mini by default in this paper), the “teacher” LLM that will generate the out-of-scope questions.
-We prefer but do not require LLMq to be larger and stronger than the LLMr under evaluation.
-For each document $d$, we call LLMq to generate a list of factual claims, then iteratively replace subsets of claims with hallucinated claims while maintaining semantic similarity to $d$ using guided hallucination, see Algorithm 1.
-We also generate in-scope questions for each document $d$ by calling LLMq directly.
-3.2 Evaluation
-We first assessed the quality of ScopeQA by sampling a subset of questions and manually verifying whether they were indeed out-of-scope or not (Section 3.3).
-Next, we evaluated how well various LLMr models could defuse the out-of-scope questions in ScopeQA.
+Thus, we select news articles with more than 150 words and, for each, sequentially extract sentences from the beginning until the word count exceeds 300.
+Finally, we collected 2,000 news and sampled a subset for human annotation.
+We represent the human-annotated data as Gold and the remaining data as Silver.
+Table 2 presents the dataset statistics.
+3.2. Question Generation
+We introduce our guided hallucination method, designed to generate out-of-scope questions from a given document.
+The process consists of three main steps: claim extraction, hallucination injection, and question generation.
+3.2.1. Claim Extraction
+The first step involves extracting claims from the document $d$ to serve as the basis for generating out-of-scope questions.
+Since our goal is to generate questions that appear related to $d$ but are actually unanswerable using its content, we start by obtaining a structured set of claims from the document.
+To achieve this, we use a LLM denoted as LLMq, to extract a set of $n$ claims $c_{1},c_{2},…,c_{n}$ from $d$.
+These claims represent key factual statements or assertions made in the document.
+We extract the claims in batches to minimize the number of LLM calls and improve computational efficiency.
+Once the claims are extracted, we partition them into disjoint subsets $S_{1},S_{2},…,S_{k}$.
+A simple partitioning scheme is used, such as:
+$S_{j}=\{c_{i}\mid i\mod 3=j-1\},\quad\text{for }j=1,2,3.$
+This ensures that claims are distributed evenly among the subsets, facilitating an efficient iterative process in the next step.
+Partitioning the claims allows us to selectively manipulate certain claims while preserving the overall document structure.
+3.2.2. Hallucination Injection
+The second step is to generate hallucinated claims that are similar to the factual claims generated in Section 3.2.1 but contain altered or missing information.
+The process follows an iterative masking and recovery approach over multiple rounds (e.g.
+3).
+For each round:
+  • For all claims in one subset $S_{j}$, we replace them with placeholders “(missing)”.
+  • The modified claims list (with missing claims) is then provided to LLMq without access to $d$. The model is prompted to recover the missing claims based on the remaining claims in the list.
+  • The recovered claims are inserted back into the claims list, replacing the placeholders.
+  • This process is repeated for all subsets $S_{1},…,S_{k}$ across several rounds.
+Finally, we discard claims that are explicitly supported by $d$. This ensures that only truly out-of-scope claims remain in our dataset.
+3.2.3. Question Generation
+The final step converts the hallucinated claims into out-of-scope questions while ensuring they cannot be answered using the document.
+Specifically, for each hallucinated claim, LLMq generates a concise and precise question focusing on a key aspect of the claim.
+The question is designed to appear related to the document while being unanswerable based on its content.
+Once the questions are generated, we further filter them by ensuring they are truly unanswerable.
+This is achieved by prompting LLMq to verify whether the generated questions can be answered using $d$.
+Any question that remains answerable in the context of $d$ is discarded.
+In addition to generating out-of-scope questions, we also prompt LLMq to generate in-scope questions for comparison.
+The complete question generation process, along with its prompts, is outlined in Algorithm 1.
+Group Name Cohen’s Kappa Annotator Acc Group’s Agree Acc Ground Truth Acc
+Out-of-Scope defusion Out-of-Scope Defusion Out-of-Scope Defusion Out-of-Scope Defusion
+1   — 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
+  — 87.50 94.59
+2   — 0.7508 0.8187 93.06 97.14 98.41 96.77
+  — 91.67 90.00
+3   — 0.8615 0.8333 91.67 93.75 98.51 100.00
+  — 98.61 100.00
+Table 3. Agreement scores and the performance of our method on the ELOQ-Gold.
+3.3. Evaluation
+We first assessed the quality of ELOQ by sampling a subset of questions (ELOQ-Gold) and manually verifying whether they were truly out-of-scope (Section 3.4).
+Next, we evaluated how well various LLMr models could defuse the out-of-scope questions in ELOQ.
 Given the large number of out-of-scope questions, manual verification of LLMr responses is not feasible.
 Therefore, we proposed AutoDefuseEval using GPT-4o-mini to automatically detect whether a response successfully defuses the question.
-To validate the effectiveness of this defusion detection method, we compared its performance with human annotations on a sampled dataset (Section 3.3).
+To validate AutoDefuseEval, we compared its performance with human annotations (ELOQ-Gold) on a sampled dataset (Section 3.4).
 Formally, we define our tasks as follows:
-Question generation:
+Question generation::
 What is the quality of out-of-scope and in-scope questions generated by Algorithm 1? (Section 3.4)
-Defusion detection:
-How accurately can AutoDefuseEval detect if an LLMr response to a out-of-scope question defuses the confusion? (Section 3.3)
-Out-of-scope response:
-How frequently does an LLMr, when responding to an out-of-scope question using different prompting methods, generate a response that successfully defuses the confusion? (Section 4.1)
-Out-of-scope detection:
+Defusion detection::
+How accurately can AutoDefuseEval detect whether an LLMr response to a out-of-scope question defuses the confusion? (Section 3.4)
+Out-of-scope response::
+How often does an LLMr successfully defuse an out-of-scope question using different prompting methods? (Section 4.1)
+Out-of-scope detection::
 How accurately can LLMs identify which context-grounded questions are out-of-scope and require special handling? (Section 4.2)
-Appendix A has the steps we run for evaluating the above tasks.
+Algorithm 2 has the steps we run for evaluating the above tasks.
 We run steps 1, 2, and 3 for “out-of-scope detection” and 1, 4, 5, 6, and 7 for “out-of-scope response”.
-Following the self-consistency method Wang et al.
-(2023a), we perform multiple LLM calls in steps  2 and  5 of Algorithm A, taking the majority vote to determine the final label.
+Following the self-consistency method (Wang et al., 2023b), we perform multiple LLM calls in steps  2 and  5 of Algorithm 2, taking the majority vote to determine the final label.
+3.4. Human Annotation
+Figure 1. Confusion matrix of out-of-scope and defusion on ELOQ-Gold.
+Following the methodology from (Cochran, 1977), we sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators.
+As shown in Table 3, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is out-of-scope and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency.
+Two additional annotators manually resolved all disagreements to establish ground truth labels.
+All annotators were computer science graduate students recruited via email and incentivized with complimentary meals.
+They were trained by first reviewing the annotation guidelines and then labeling a small dataset previously annotated by the authors.
+The annotation guidelines are released along with the data and code.
+We used two metrics to assess our method: “Annotator Acc” and “Group’s Agree Acc.” “Annotator Acc” treats each annotator’s labels as ground truth, computing the accuracy of the generated labels.
+“Group’s Agree Acc” uses the instances where both annotators within the same group agreed, treating this consensus as the ground truth.
+This method yielded higher accuracy than “Annotator Acc,” leading us to resolve disagreements with two additional annotators, ultimately establishing the final labeled data as ground truth.
+As reflected in “Ground Truth Acc,” in Table  3, our method achieved 94.91% accuracy in generating out-of-scope and in-scope questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion.
+As shown in Figure 1, the confusion matrix on the left aligns closely with the ground truth, with minimal errors (5 false negatives and 6 false positives).
+The defusion matrix on the right reflects a similarly high accuracy, with only one error for false negatives and false positives.
+Model Name Quantization Knowledge Cutoff Date
+Llama 3.2 3B Instruct Turbo FP8 2023/12
+Llama 3.1 8B Instruct Turbo FP8 2023/12
+Llama 3.1 70B Instruct Turbo FP8 2023/12
+Llama 3.3 70B Instruct Turbo FP8 2023/12
+Mistral (7B) Instruct v0.3 FP16 2024/5
+gpt-3.5-turbo   — 2021/9
+Table 4. Different LLMr being evaluated. In ELOQ-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cutoff date of 5/22/2024.
+4. Experimental Results
 Prompt Model business entm food music news politics science sport tech travel Avg Std Dev
@@
 Llama 3.3 70B 93.52 91.24 90.55 92.57 90.77 92.12 89.12 $\textbf{95.09}^{\star}$ 93.85 90.90 91.97 1.71
-Table 2: Evaluation of LLMs’ accuracy in defusing out-of-scope questions from the ScopeQA-Silver dataset across diverse news topics.
-For each LLM, the underscored value and stared () value are the minimum and maximum values, respectively, across all the topics, within the same prompting method.
+Table 5.
+Evaluation of LLMs’ accuracy in defusing out-of-scope questions from the ELOQ-Silver dataset across diverse news topics.
+For each LLM, the underscored value and stared () value are the minimum and maximum values, respectively, across all the topics within the same prompting method.
 Bold values are the maximum values for each topic across all the LLMs within the same prompting method.
-3.3 Human Annotation
-To evaluate “question generation” and “defusion detection” (Section 3.2), we recruited six computer science graduate students for manual labeling (Appendix B).
-As reflected in “Ground Truth Acc,” of Appendix B, Table  4, our method achieved 94.91% accuracy in generating out-of-scope and in-scope questions, while our proposed AutoDefusionEval method achieved 98.23% accuracy in evaluating defusion.
-As shown in Appendix B, Figure 2, the confusion matrix on the left strongly agrees with the ground truth, with minimal errors (5 false negatives and 6 false positives).
-The defusion matrix on the right reflects a similarly high accuracy, with only one error for false negatives and false positives.
-4 Experimental Results
-4.1 Out-of-Scope Response
-We evaluated open-source LLMs (Appendix 5) on their ability to defuse out-of-scope questions with three prompting methods (Appendix F.6,  F.7,  F.8).
+“entm” is the abbreviation of “entertainment”.
+4.1. Out-of-Scope Response
+In this section, we examine LLMs’ ability to defuse out-of-scope questions.
+The selected LLMs are listed in Table 4.
+Most of these models have a knowledge cutoff date of December 2023, which predates the publication of the news articles in ELOQ.
+As a result, they must generate responses based solely on their understanding of the provided news content.
+As shown in Table 5, we evaluated LLMs on their ability to defuse out-of-scope questions with three prompting methods (Appendix B.6,  B.7,  B.8).
 As shown in Table 5, a) Two-shot prompt boosts all evaluated LLMs’ accuracy substantially, especially GPT-3.5, likely due to the examples clarifying how to respond to out-of-scope questions.
-In contrast, Zero-shot-CoT Kojima et al.
-(2022) merely invites the LLM to “reason step by step” beats Two-shot on three out of six LLMs.
-b) Larger models (70B) benefit more from Zero-shot-CoT than from Two-shot prompting, achieving accuracy gains of 5.07% to 9.98%, while smaller models (3B to 8B) experience only marginal changes, ranging from -3.68% to 1.67%.
-This difference likely stems from larger models’ ability to better utilize their advanced reasoning capabilities and extensive knowledge; c) Most LLMs perform worst on the “science” topic and best on “sport”.
+In contrast, Zero-shot-CoT (Kojima et al., 2022) merely invites the LLM to “reason step by step” beats Two-shot on three out of six LLMs.
+b) Larger models (70B) benefit more from Zero-shot-CoT than from Two-shot prompting, achieving accuracy gains of 5.07% to 9.98%, whereas smaller models (3B to 8B) show only marginal changes, ranging from -3.68% to 1.67%.
+This difference likely stems from larger models’ ability to better utilize their reasoning capabilities and extensive knowledge; c) Most LLMs perform worst on the “science” topic and best on “sport”.
 This is expected, as scientific content tends to contain more implicit premises assumed by the author rather than explicitly stated facts, thus requiring more domain knowledge.
-In contrast, sports content is simpler and unambiguous, and the facts in the document are often clearly stated.
-4.2 Out-of-Scope Detection
-By default, we use GPT-4o-mini as our LLMq for “out-of-scope detection” task.
-In this section, we study the impact of different LLMs and methods.
-As illustrated in Figure 1, “Direct Generate” prompts the LLM to do the detection (Appendix F.3); “Unused-token Classification” represents a classifier trained on ScopeQA-Silver using the embedding of an appended unused-token (Appendix H) as a feature.
-Finally, “SBERT NLI” represents a classifier trained on ScopeQA-Silver using embeddings from a sentence-transformer model (Appendix H.3).
-Among these methods, GPT-4o-mini performs best with the “Direct Generate” approach.
-Notably, ScopeQA proves its utility by enabling even smaller models, like the Llama 3.2 3B-based classifier, to achieve accuracy comparable to much larger models, such as Llama 3.1 70B and GPT-4o-mini.
-To minimize interference with the LLM’s generation process, we append an unused-token, leaving its potential impact for future investigation.
-Figure 1: Evaluation of out-of-scope detection on ScopeQA-Gold.
+In contrast, sports content is simpler and unambiguous, with facts often clearly stated in the document.
+4.2. Out-of-Scope Detection
+The main challenge in preventing an LLM from hallucinating or generating incorrect answers to out-of-scope questions is detecting such scenarios.
+Once we can accurately identify these cases, we can handle them easily by using a prompt and in-context examples specifically designed for out-of-scope questions.
+We use two methods for out-of-scope detection.
+First is a prompt-based method where we simply prompt the LLM to do such detection (Appendix B.3), denoted as “Direct Generate” in Figure 2.
+Second is a classification method where we train a binary classifier on ELOQ-Silver using the embeddings of an appended unused-token as features, denoted as “Unused-token Classification” in Figure 2.
+Figure 2. Evaluation of out-of-scope detection on ELOQ-Gold.
+4.2.1. Unused-token Classification
+The out-of-scope detection problem can be framed simply as a binary classification task, where we determine if a given query $q$ is out-of-scope or not for a given document $d$.
+The binary label is $y\in\{0,1\}$ where $y=1$ indicates the query is relevant to the document and $y=0$ indicates the query is out-of-scope.
+The objective is to learn a classification function: $f:(q,d)\mapsto p(y=1\mid q,d)$, which estimates the likelihood of the query-document pair to be in-scope.
+We would like to test whether the LLM’s internal features contain enough information to classify a query-document pair effectively to simulate the effect that the LLM “realizes” the out-of-scope scenario before generating a response.
+To achieve this, we use the sequence representation of the entire input – denoted as $x=\text{concat}(I,q,d,)$ where $I$ is the instruction prompt provided to the LLM (e.g.
+“Can this document answer the question?”).
+We extract the sequence representation by appending a reserved unused-token to the end of the input sequence and extracting its representation from the final hidden state of the last transformer layer.
+The rationale behind using unused-token is that, since LLM developers reserve these tokens for potential future use in specialized training tasks, they are untrained and uninfluenced by any prior data.
+Thus, the unused-token does not have any predefined meanings and can be viewed as a blank buffer, aggregating interactions among all the input tokens and capturing an uninfluenced representation of the sequence.
+We also tried using the <eos> token, which yielded similar results.
+The classifier is a two-layer MLP that takes the unused-token representation as input and is optimized against the binary label $y$.
+This approach is commonly known as “probing” (Ye et al., 2024), where the LLM weights are frozen, and the classifier can be viewed as an external probe to assess the information contained in the model’s internal representations.
+The performance reported in Figure 2 was tested on ELOQ-Gold, the human-labeled data.
+For more details about the training, please refer to Appendix D.
+We also provide a “Sentence BERT NLI” baseline where we adopt NLI style training using the off-the-shelf  [^4] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 model from HuggingFace.
+Specifically, we follow (Reimers and Gurevych, 2019) to extract document and question [CLS] embeddings $e_{d}$ and $e_{q}$ to use $[e_{q};e_{q};|e_{d}-e_{q}|]$ as the input features.
+4.2.2. Results Discussion
+For the “Direct Generation” approach, we can observe a clear positive correlation between model size and accuracy, aligning with the general trend of increased capability with larger models.
+However, due to the proprietary nature of OpenAI’s models and computational constraints, we are only able to obtain hidden states to train the binary classifier for three LLMs: Llama 3.2 3B, Mistral 7B, and Llama 3.1 8B.
+Notably, the classifiers consistently outperform the “Direct Generation” method, indicating that the LLM’s internal representations encode the necessary information for out-of-scope detection, yet the model may fail when directly prompted.
+Our dataset, ELOQ, demonstrates its usefulness by enabling even smaller models – such as Llama 3.2 3B-based classifier – to achieve accuracy comparable to significantly larger models that rely on direct generation, including Llama 3.1 70B and GPT-4o-mini.
+Moreover, as shown in Figure 2, all models and methods outperform the “SBERT NLI” baseline, suggesting that out-of-scope detection is non-trivial and requires a deeper and more complex understanding of linguistic relationships that larger models capture during pretraining and fine-tuning.
+Lastly, in our exploratory experiments, we trained a classifier on GPT-2, but its accuracy is no better than random guessing, further highlighting the importance of model capacity in leveraging internal representations for classification.
+Model Recall@1 Recall@5 Recall@10 MRR
+BM25 (Pérez-Iglesias et al., 2009) 0.4637 0.6880 0.7487 0.5653
+SBERT (Reimers and Gurevych, 2019) 0.4685 0.7058 0.7858 0.5773
+BGE (Xiao et al., 2023) 0.5326 0.7604 0.8267 0.6352
+Stella (Zhang and FulongWang, 2024) 0.4875 0.7376 0.8128 0.5994
+Linq (Kim et al., 2024) 0.5603 0.7890 0.8492 0.6618
+Table 6. Retrieval performance of out-of-scope questions on ELOQ
+4.3. Semantic Relevance
+To demonstrate that the generated out-of-scope questions are indeed semantically similar to the document they are generated from, we conduct retrieval experiments using all out-of-scope questions against our document corpus (2000 document).
+We employ the following retrieval models: BM25  [^5] https://github.com/castorini/pyserini, SBERT  [^6] https://huggingface.co/sentence-transformers/all-mpnet-base-v2, bge-large-en-v1.5  [^7] https://huggingface.co/BAAI/bge-large-en-v1.5, stella_en_1.5B_v5888https://huggingface.co/NovaSearch/stella_en_1.5B_v5, and Linq-Embed-Mistral [^9] https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral.
+Among these models, BM25 and SBERT are well-known retrievers that serve as established baselines, bge-large-en-v1.5 was accepted in SIGIR 2024 resource track as a high-performance embedding model, stella_en_1.5B_v5 ranks #1 on MTEB’s English v1 dataset (Muennighoff et al., 2022), and Linq-Embed-Mistral leads multiple benchmarks, including BEIR (Kamalloo et al., 2023) and MTEB’s English v2 dataset at the time of writing.
+As shown in Table 6, the Recall@1 values across all models suggest that for approximately half of the out-of-scope questions, the top-retrieved document is the same one from which the question was generated.
+When we relax this condition to Recall@10, around 80% of out-of-scope questions retrieve their associated document within the top 10 results.
+This demonstrates the strong semantic connection between out-of-scope questions and their source documents despite the fact that the document lacks the information needed to answer the question.
 Module Accuracy
@@
 + self-consistency (m=9) 98.23
-Table 3: Evaluation of different prompts for “defusion detection” on ScopeQA-Gold.
-4.3 Ablation Study
+Table 7. Evaluation of different prompts for “defusion detection” on ELOQ-Gold.
+4.4. Ablation Study
 We examined the impact of various prompts on “defusion detection” task.
@@
 However, further increasing $m$ did not lead to additional improvements.
-Since ScopeQA-Gold is a smaller dataset, using $m=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger ScopeQA-Silver dataset.
-Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix G).
-5 Conclusion
-We introduced ScopeQA, a benchmark for evaluating LLMs’ ability to defuse out-of-scope questions in RAG.
-Human annotation validates that ScopeQA achieves 94.91% accuracy in distinguishing out-of-scope questions, while our AutoDefusionEval achieves 98.23% accuracy in detecting effective defusion responses.
-We further demonstrated ScopeQA’s utility by training out-of-scope detectors.
-For future work, we plan to investigate the impact of appending an unused-token on the generation process.
-Additionally, we aim to explore instruction fine-tuning by training LLMs on ScopeQA and evaluating their performance.
-We also intend to apply for OpenAI funding to leverage more powerful models, such as GPT-4o, to enhance data quality.
-6 Acknowledgement
-We would like to express our sincere gratitude to Xiaoxiao Shang, Ethan Lin, Wei Mo, Zhan Shi, Xinze Ye, and Haisong Wang for their invaluable contributions to data labeling.
-7 Limitations
-In this paper, we evaluated LLMs’ abilities to defuse out-of-scope questions using three prompting methods: basic prompt, Two-shot prompt, and Zero-shot-CoT prompt.
-These methods were chosen because LLMs should ideally handle out-of-scope questions with straightforward prompts.
-However, we did not assess the performance of LLMs fine-tuned on ScopeQA-Silver due to computational resource constraints.
-Additionally, we empirically appended an unused-token to the end of the prompt for classification while minimizing its influence on LLMs’ generation processes.
-The effect of appending such tokens on generation remains unexplored and is left for future work.
-Finally, budget constraints prevented us from using more powerful models, such as GPT-4o or GPT-o1, which could theoretically produce higher-quality datasets.
-Future users can adopt our framework with more advanced models and alternative data sources to generate even higher-quality data tailored to their needs.
-Appendix A Algorithm
-Algorithm 2 Evaluation steps
-1:Call LLMr on $(d,q)$ with RAG prompt and get its response $r$ (Appendix F.6,  F.7,  F.8)
-2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is out-of-scope given $d$ (Appendix F.3)
-3:Aggregate $m$ predictions by majority vote
-4:if $q$ is generated as out-of-scope then
-5:     Call LLMq $m$ times on $(d,q,r)$ to check and explain if $r$ defused the confusion (Appendix F.4)
-6:     Aggregate $m$ predictions by majority vote
-7:end if
-Appendix B Human Annotation
-Figure 2: Confusion matrix of out-of-scope and defusion on ScopeQA-Gold.
-Group Name Cohen’s Kappa Annotator Acc Group’s Agree Acc Ground Truth Acc
-Out-of-Scope defusion Out-of-Scope Defusion Out-of-Scope Defusion Out-of-Scope Defusion
-1   — 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
-  — 87.50 94.59
-2   — 0.7508 0.8187 93.06 97.14 98.41 96.77
-  — 91.67 90.00
-3   — 0.8615 0.8333 91.67 93.75 98.51 100.00
-  — 98.61 100.00
-Table 4: Agreement scores and the performance of our method on the ScopeQA-Gold.
-Following the methodology from Cochran (1977), we sampled 216 questions, evenly divided into three groups, ensuring that each question was annotated by two different annotators.
-As shown in Table 4, Cohen’s Kappa values exceeded 0.75 for confusion labels (where “Yes” indicates a question is out-of-scope and “No” indicates otherwise), and surpassed 0.81 for defusion labels (where “Yes” indicates the LLMr’s response defuses the question, and “No” indicates it does not), signifying substantial to near-perfect inter-annotator agreement and annotation consistency.
-Two more annotators manually resolved all disagreements to establish ground truth labels.
-All the annotators are computer science graduate students recruited via email and incentivized with complimentary meals.
-All annotators were trained by first reviewing the annotation guidelines and then labeling a small dataset previously labeled by the authors.
-The annotation guideline and annotation tool will be released along with the data and code upon acceptance.
-We used two metrics to assess our method: “Annotator Acc” and “Group’s Agree Acc.” “Annotator Acc” treats each annotator’s labels as ground truth, computing the accuracy of our method’s generated labels.
-“Group’s Agree Acc” uses the instances where both annotators within the same group agreed, treating this consensus as the ground truth.
-This method demonstrated better accuracies than “Annotator Acc,” leading us to resolve disagreements with two additional annotators, ultimately establishing the final labeled data as ground truth.
-Appendix C Different LLMr
-Model Name Quantization Knowledge Cut-off Date
-Llama 3.2 3B Instruct Turbo FP8 2023/12
-Llama 3.1 8B Instruct Turbo FP8 2023/12
-Llama 3.1 70B Instruct Turbo FP8 2023/12
-Llama 3.3 70B Instruct Turbo FP8 2023/12
-Mistral (7B) Instruct v0.3 FP16 2024/5
-gpt-3.5-turbo   — 2021/9
-Table 5: Different LLMr to be evaluated. In ScopeQA-Silver, 66 news articles are published before Mistral 7B v0.3’s knowledge cut-off date 5/22/2024.
-Appendix D Out-of-scope Example
+Since ELOQ-Gold is a smaller dataset, using $m=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger ELOQ-Silver dataset.
+Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix C).
+5. Conclusion and Future Works
+We introduced ELOQ, a benchmark for enhancing LLM detection of out-of-scope questions for a given document.
+Human annotation validates that ELOQ achieves 94.91% accuracy in generating out-of-scope and in-scope questions, while our AutoDefusionEval achieves 98.23% accuracy in detecting effective defusion responses.
+We further demonstrated ELOQ’s utility by training out-of-scope detectors.
+For future work, we plan to apply for OpenAI funding to utilize more powerful models, such as o1, to improve both data quality and scale.
+Additionally, we aim to explore post-training techniques to fine-tune LLMs on ELOQ, ensuring they align better with human preferences by clarifying the confusing aspects of a question rather than generating hallucinated answers.
+Appendix A Out-of-scope Example
 Document:
@@
 Question: How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?
-Explanation: This question is out-of-scope as it asks about "Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft" which was never discussed in the document.
-Appendix E Statistics of ScopeQA
-Attr Gold Silver
-# of documents 45 1955
-# of in-scope questions 103 8,949
-# of out-of-scope questions 113 10,105
-average words per document 280 276
-average words per in-scope question 16 16
-average words per out-of-scope question 16 16
-Table 6: Statistics of ScopeQA. Gold is sampled from crawled 2000 news and annotated by annotators (Appendix B).
-Appendix F Prompt
-F.1 Extract Claims Prompt
+Explanation: This question is out-of-scope as it asks about ”Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft” which was never discussed in the document.
+Appendix B Prompt
+B.1. Extract Claims Prompt
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@
 Document:
-"""{document}"""
+”””{document}”””
 {num_fact} most important facts:
-F.2 Recover Missing Claims Prompt
+B.2. Recover Missing Claims Prompt
 Read the document below with a list of {num_fact} facts it contains.
@@
 Document:
-"""document"""
+”””document”””
 Completed list of facts:
-F.3 Out-of-scope Judgement Prompt
+B.3. Out-of-scope Judgement Prompt
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@
 Document:
-"""{document}"""
+”””{document}”””
 Question:
@@
 End your response with ’The answer is: Yes or No.’:
-F.4 Response Defusion Judgement Prompt
+B.4. Response Defusion Judgement Prompt
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions. Read the document and answer the question based on the document.
 Document:
-"""{example_document}"""
+”””{example_document}”””
 Question:
@@
 {example_LLM_responses}
-{Defusion Instruction (Appendix  F.5)}
+{Defusion Instruction (Appendix  B.5)}
 {example_defuse_reasoning}
@@
 Document:
-"""{document}"""
+”””{document}”””
 Question:
@@
 {LLM_response}
-{Defusion Instruction (Appendix  F.5)}
-F.5 Defusion Instruction
+{Defusion Instruction (Appendix  B.5)}
+B.5. Defusion Instruction
 The question presented is confusing and unanswerable based on the document provided.
@@
 Note that if the response attempts to answer the question at any point, like answering ’Yes’ to agree with the question or ’No’ to disagree with the question, you should immediately conclude with ‘The answer is: No.’ even if it later acknowledges the question’s confusion:
-F.6 RAG Basic
+B.6. RAG Basic
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@
 Document:
-"""{document}"""
+”””{document}”””
 Question:
@@
 Answer:
-F.7 RAG Two-shot
+B.7. RAG Two-shot
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@
 Document:
-"""{example_document}"""
+”””{example_document}”””
 Question:
@@
 Document:
-"""{document}"""
+”””{document}”””
 Question:
@@
 Answer:
-F.8 RAG Zero-shot-CoT
+B.8. RAG Zero-shot-CoT
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@
 Document:
-"""{document}"""
+”””{document}”””
 Question:
@@
 Answer:
-F.9 In-scope Question Generation Prompt
+B.9. In-scope Question Generation Prompt
 You will be provided with a document delimited by triple quotes.
@@
 Document:
-"""{document}"""
+”””{document}”””
 Questions:
-F.10 Out-of-scope Question Generation Prompt
+B.10. Out-of-scope Question Generation Prompt
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@
 Document:
-"""{document}"""
+”””{document}”””
 hallucinated facts:
@@
 Questions:
-F.11 Remove Claims Prompt
+B.11. Remove Claims Prompt
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@
 Document:
-"""{example_document}"""
+”””{example_document}”””
 {num_true_fact} ground-truth facts:
@@
 Document:
-"""{document}"""
+”””{document}”””
 {num_true_fact} ground-truth facts
@@
 Remaining hallucinated facts:
-Appendix G Expenses
-Generating the ScopeQA dataset using GPT-4o-mini costs approximately \$80.
+Appendix C Expenses
+Generating the ELOQ dataset using GPT-4o-mini costs approximately \$80.
 For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/.
 Excluding the dataset generation, all other experiments cost around \$100.
-Appendix H Embedding Classifier
-H.1 Implementation Details
-We train the classifier on ScopeQA-Silver, which is split as: 80% for training, 10% for evaluation, and 10% for testing. After training, we test the classifier on ScopeQA-Gold data.
-We append the unused-token ID at the end of the tokenized input sequence.
-After the LLM processes the input, we extract the unused-token’s hidden state from the model’s last hidden layer output.
-Since this token is not part of the LLM’s training and does not carry predefined meanings, it serves as a placeholder to aggregate information from the model’s attention over all preceding tokens in the sequence.
-As the final token in the input, its embedding encapsulates the model’s representation of the entire input sequence, reflecting the information inferred from the input sequence.
+Appendix D Embedding Classifier Implementation Details
+We train the classifier on ELOQ-Silver, which is split as: 80% for training, 10% for evaluation, and 10% for testing. After training, we test the classifier on ELOQ-Gold data.
+For the Llama family, we use <|reserved_special_token_0|> as the unused-token.
+This token is not trained and is reserved for future use cases such that the users do not need to resize the vocabulary.
+For Mistral-7B-v0.3, we use [control_555] as the unused-token because Mistral’s official documentation indicated that they reserve 768 control tokens for future use, and none of them are used during training.
+555 is a random choice.
 The unused-token’s hidden state is used as input features for a 2-layer MLP, which is optimized with a BCE loss.
@@
 We train each classifier for 10 epochs with batch size 8.
-We save the best validation checkpoint and report its performance on ScopeQA-Gold.
-H.2 Unused-Token
-For the Llama family, we use the <|reserved_special_token_0|> as the unused-token.
-This token is not trained and is reserved for future use cases such that the users do not need to resize the vocabulary.
-For Mistral-7B-v0.3, we use [control_555] as the unused-token because Mistral’s official documentation indicated that they reserve 768 control tokens for future use, and none of them are used during training.
-555 is a random choice.
-H.3 Sentence BERT NLI baseline
-To create a baseline classification performance, we adapt NLI style training using an off-the-shelf sentence-transformers/all-mpnet-base-v2 model from HuggingFace.
-Specifically, we follow Reimers and Gurevych (2019) to extract document and question [CLS] embeddings $u$ and $v$ to use $[u;v;|u-v|]$ as the input feature for the 2-layer MLP to classify.
+The best validation checkpoint is saved and its performance is reported on ELOQ-Gold.
+Acknowledgements.
+We would like to express our sincere gratitude to Xiaoxiao Shang, Ethan Lin, Wei Mo, Zhan Shi, Xinze Ye, and Haisong Wang for their invaluable contributions to data labeling.

### v4

@@
 Abstract.
-Large Language Models (LLMs) are widely used in Conversational AI systems to generate responses to user inquiries.
-However, many natural questions lack well-defined answers.
-While existing studies primarily focus on question types such as false premises, they often overlook out-of-scope questions, where the provided document is semantically highly similar to the query but does not contain the required answer.
-In this paper, we propose a guided hallucination-based method to efficiently generate a diverse set of out-of-scope questions from a given document corpus.
-We then evaluate multiple LLMs based on their effectiveness in confusion detection and appropriate response generation.
-Furthermore, we introduce an improved method for detecting such out-of-scope questions, enhancing the reliability of LLM-based question-answering systems.
-Large Language Models, Question Answering, Out-of-Scope Question
-*Both authors contributed equally to this work.
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processing
+Retrieval-augmented generation (RAG) has become integral to large language models (LLMs), particularly for conversational AI systems where user questions may reference knowledge beyond the LLMs’ training cutoff.
+However, many natural user questions lack well-defined answers, either due to limited domain knowledge or because the retrieval system returns documents that are relevant in appearance but uninformative in content.
+In such cases, LLMs often produce hallucinated answers without flagging them.
+While recent work has largely focused on questions with false premises, we study out-of-scope questions, where the retrieved document appears semantically similar to the question but lacks the necessary information to answer it.
+In this paper, we propose a guided hallucination-based approach ELOQ  [^1] https://github.com/zhiyuanpeng/ELOQ.git, to automatically generate a diverse set of out-of-scope questions from post-cutoff documents, followed by human verification to ensure quality.
+We use this dataset to evaluate several LLMs on their ability to detect out-of-scope questions and generate appropriate responses.
+Finally, we introduce an improved detection method that enhances the reliability of LLM-based question-answering systems in handling out-of-scope questions.
+Large Language Models, Question Answering, Out-of-Scope Question, Retrieval Augmented Generation
+ccs: Computing methodologies Natural language processingccs: Information systems Question answering
 1. Introduction
 Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents (Lewis et al., 2020; Guu et al., 2020; Gao et al., 2024).
-Given an inquiry, a RAG system retrieves a relevant document from a curated knowledge base and provides it to a large language model (LLM) that generates a response.
+Given an inquiry, a RAG system retrieves a relevant document from a curated knowledge base and presents it to a large language model (LLM), which generates a response.
 RAG performance is evaluated on dimensions such as response accuracy, faithfulness, completeness, answer and context relevance, and style (Es et al., 2024; Saad-Falcon et al., 2024; Liu et al., 2024; Zheng et al., 2023).
@@
 Additionally, out-of-scope errors have been identified as the leading cause of enterprise AI assistants generating responses that appear convincing but are, in fact, incorrect (Maharaj et al., 2024).
-A well-designed RAG system should implement an exception-handling mechanism that detect confusing questions and responds by clarifying the confusion, retrieving additional documents, or seeking assistance.
-Dataset Source Category Cutoff knowledge Annotation Method
+A well-designed RAG system should implement an exception handling mechanism that detects confusing questions and responds by clarifying the confusion, retrieving additional documents, or seeking assistance.
+Dataset Source Category Cutoff Annotation Method
+SQuAD 2.0 (Rajpurkar et al., 2018) Wiki Unanswerable Human
+SQuAD2-CR (Lee et al., 2020) Wiki Unanswerable Human
+BoolQ3L (Sulem et al., 2022) Wiki Unanswerable Algoritm+Human
+CAsT-answerability (Lajewska and Balog, 2024) Wiki Unanswerable Algoritm+Human
 AMBIGQA (Min et al., 2020a) Wiki Multi-answers Human
@@
 FLUB (Li et al., 2024a) Baidu Tieba Cunning Texts Human
-ELOQ News Out-of-scope Human+LLM
-Table 1. ELOQ vs Existing Datasets
+ELOQ News Out-of-scope LLM+Human
+Table 1. ELOQ vs Existing Datasets. “Cutoff” represents whether the data is beyond the LLMs’ knowledge or not.
 Enabling “deconfusion” in a RAG system involves three challenges: (1) creating a diverse dataset of confusing questions, (2) training a classifier to detect confusion, and (3) developing a response generator for each confusion type.
-As shown in Table 1, much prior work focuses on false-premise questions that contradict pretraining knowledge, but often neglects out-of-scope queries for which no answers can be found in retrieved documents.
-Out-of-scope questions can mislead LLMs into generating hallucinated answers that appear correct, but are incorrect, which represents the most dangerous type of error (Maharaj et al., 2024).
-Our focus is to mitigate this type of confusion.
-This paper proposes a framework for automatically generating out-of-scope questions and evaluates its effectiveness in training classifiers to detect out-of-scope situations.
-While queries in MS MARCO (Nguyen et al., 2016) may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cutoff.
+As shown in Table 1, prior studies create confusing questions with different categories, such as multi-answer (Min et al., 2020a), false premise (Yu et al., 2023; Hu et al., 2023; Zhu et al., 2024; Yuan et al., 2024), and cunning texts (Li et al., 2024a).
+For instance, “How many sacks does clay matthews have in his career” is a multi-answer question as “clay matthews” may refer to “Clay Matthews Jr.” or “Clay Matthews III”, “Where in Liverpool did John Lennon die” is a false premise question as John Lennon died in New York city and “What should I do if I forget which ATM machine I deposited my money in?” is a cunning text.
+In contrast to these questions, out-of-scope questions look relevant to the document, but there is no answer within the document and thus can mislead LLMs into generating hallucinated answers that seem correct but are actually incorrect, representing the most dangerous type of error (Maharaj et al., 2024).
+For instance, “How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?” is an out-of-scope question as it asks about “Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft” which was never discussed in the paired news (Appendix D).
+Our work focuses on mitigating this type of confusion.
+Existing studies (Rajpurkar et al., 2018; Lee et al., 2020; Sulem et al., 2022; Lajewska and Balog, 2024) on unanswerable questions of which out-of-scope questions are a subset, have typically created the dataset either through manual annotation by trained human annotators (Lee et al., 2020; Hu et al., 2023) or via simple automatic generation algorithms, as seen in BoolQ3L (Sulem et al., 2022) and CAsT-answerability (Lajewska and Balog, 2024).
+These methods often lack a balance between efficiency and quality, as human-annotated questions are of high quality but low efficiency, and automatically generated questions are of high efficiency but too simple to be distinguished to some degree.
+Recently, LLMs have demonstrated strong generation and following instruction abilities, and thus KG-FPQ (Zhu et al., 2024) and FAITH (Yuan et al., 2024) utilize LLMs to generate high-quality false premise questions, mitigating the balance of efficiency and quality issues.
+However, all these datasets are usually constrained to knowledge before the LLMs’ training cutoff date, which limits their effectiveness in evaluating real-world RAG systems that must retrieve up-to-date information likely beyond the model’s training data to answer user questions.
+To address these challenges above, we propose a framework for automatically generating out-of-scope questions and evaluate its effectiveness in training classifiers to detect out-of-scope situations.
+While questions in existing datasets such as MS MARCO (Nguyen et al., 2016), may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cutoff.
 ELOQ fills this gap by providing a diverse benchmark for out-of-scope questions, leveraging LLM-assisted generation to reduce human annotation effort.
 Our contributions are as follows:
-  • We propose and implement an efficient method for generating synthetic out-of-scope questions from a given set of documents.
-  • We demonstrate the utility of our synthetic data in training out-of-scope exception detectors.
-  • We conduct a comparative evaluation of multiple LLMs as RAG agents to measure their confusion detection accuracy.
-  • Our code  [^1] https://github.com/zhiyuanpeng/ELOQ.git and dataset  [^2] https://huggingface.co/datasets/zhiyuanpeng/ELOQ are made publicly available.
+  • We propose a framework for generating synthetic out-of-scope questions from a given corpus.
+  • We demonstrate the utility of our synthetic data in training out-of-scope question detectors.
+  • We conduct a comparative evaluation of multiple LLMs as RAG agents to measure their out-of-scope question detection accuracy.
 2. Related Work
@@
 These questions may be naturally collected from online sources (Li et al., 2024b; Yu et al., 2023), written by human annotators (Hu et al., 2023), or synthetically generated (Zhu et al., 2024; Yuan et al., 2024).
-Released benchmark datasets with confusing questions include KG-FPQ (Zhu et al., 2024), FalseQA (Hu et al., 2023), FLUB (Li et al., 2024b), and CREPE (Yu et al., 2023).
-Earlier research on ambiguous open-domain questions (Min et al., 2020b) introduced AmbigNQ, a dataset of naturally ambiguous questions.
+Earlier research on ambiguous open-domain questions introduced AmbigQA (Min et al., 2020b), a dataset of naturally ambiguous questions.
+More relevant benchmarks are unanswerable questions (Rajpurkar et al., 2018; Lee et al., 2020; Sulem et al., 2022; Lajewska and Balog, 2024) of which out-of-scope questions are a subset.
+SQuAD 2.0 (Rajpurkar et al., 2018) is created by asking crowdworkers to write unanswerable questions similar to the answerable question given a paragraph.
+SQuAD2-CR (Lee et al., 2020) tags SQuAD 2.0 (Rajpurkar et al., 2018) with answerable reasons.
+CAsT-answerability (Lajewska and Balog, 2024) extends CAsT-snippets (Łajewska and Balog, 2023) by including five randomly selected non-relevant passages to each question.
+BoolQ3L (Sulem et al., 2022) matches each selected “Yes” or “No” question of BoolQ (Clark et al., 2019) to a passage within BoolQ that has the greatest overlap with the questions in terms of nouns and verbs.
+These unanswerable datasets are sourced from Wikipedia, which is typically included in the training data of LLMs.
+Our dataset is crawled from news, allowing us to easily select news published only after the knowledge cutoff date of LLMs.
 Table 1 presents a comparison of these benchmark datasets.
 2.2. Question Generation
-(Zhu et al., 2024; Yuan et al., 2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of (subject, relation, object) using Wikipedia data, replacing the object with a similar but incorrect one, and then filling in a template to generate a false premise question.
+Zhu et al.
+(2024) and Yuan et al.
+(2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of (subject, relation, object) using Wikipedia data, replacing the object with a similar but incorrect one, and then filling in a template to generate a false premise question.
 Content-grounded (non-confusing) synthetic data generation has been studied in (Yehudai et al., 2024; Wang et al., 2023a; Lee et al., 2023).
@@
 Experiments show that fine-tuning outperforms both in-context learning and CoT reasoning (Hu et al., 2023).
-CoT reasoning leads to inconsistent improvements, whereas in-context learning improves performance as the number of examples increase (Li et al., 2024b).
+CoT reasoning leads to inconsistent improvements, whereas in-context learning improves performance as the number of examples increases (Li et al., 2024b).
 Yuan et al.
@@
 (2021) retrieves a Wikipedia page for each question and responds with the identified unverifiable premises if the page does not entail them.
+Sulem et al.
+(2022) trains a classification model based on BERT to identify unanswerable questions.
 Algorithm 1 The guided hallucination method
@@
 We developed the ELOQ data generator to evaluate and mitigate LLM confusion when responding to out-of-scope questions for a given document.
-We assume that the user has limited knowledge of the document, resulting in asking questions that appear relevant but are actually out-of-scope and cannot be answered based on the document’s content.
+We assume that the user has limited domain knowledge, resulting in asking questions that appear relevant but are actually out of scope and cannot be answered based on the document’s content.
+We also assume that, due to the limitations of the retriever itself, relevant documents without an answer to the questions are possibly retrieved and ranked at the top.
+The best answer to this kind of question is to refuse to answer it, rather than providing a fabricated response.
 In this step, given a standard prompt $p$, an out-of-scope question $q$, and a relevant document $d$, the LLMr generates a response $r$.
 Ideally, $r$ should clarify the confusing part of the question – a process we refer to as defusion (or de-confusion) – rather than attempting to answer the question directly and risking hallucination.
-Algorithm 2 Evaluation steps
-1:Call LLMr on $(d,q)$ with RAG prompt and get its response $r$ (Appendix B.6,  B.7,  B.8)
-2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is out-of-scope given $d$ (Appendix B.3)
-3:Aggregate $m$ predictions by majority vote
-4:if $q$ is generated as out-of-scope then
-5:     Call LLMq $m$ times on $(d,q,r)$ to check and explain if $r$ defused the confusion (Appendix B.4)
-6:     Aggregate $m$ predictions by majority vote
-7:end if
-Attr Gold Silver
-# of documents 45 1955
-# of in-scope questions 103 8,949
-# of out-of-scope questions 113 10,105
-average words per document 280 276
-average words per in-scope question 16 16
-average words per out-of-scope question 16 16
-Table 2. Statistics of ELOQ. Gold is sampled from crawled 2000 news and annotated by annotators (Section 3.4).
 3.1. Data Collection
@@
 Instead, they must generate responses grounded in their understanding of the news content.
-Therefore, using the Newscatcher (Bugara et al., 2020), we collected 200 news articles published between Jan 1, 2024, and September 30, 2024, for each topic333sport, business, science, food, politics, travel, entertainment, music, news, tech.
-We require each document to be short enough to fit in the LLM prompt along with instructions and examples but long enough to make at least 4 to 10 separate claims (i.e.
-a few paragraphs long).
+Therefore, using the Newscatcher (Bugara et al., 2020), we collected 200 news articles published between Jan 1, 2024, and September 30, 2024, for each topic222sport, business, science, food, politics, travel, entertainment, music, news, tech.
+We require each document to be concise enough to fit within the LLM prompt, along with instructions and examples, but sufficiently long to make at least 4 to 10 separate claims (i.e., a few paragraphs in length).
 Thus, we select news articles with more than 150 words and, for each, sequentially extract sentences from the beginning until the word count exceeds 300.
@@
 We represent the human-annotated data as Gold and the remaining data as Silver.
-Table 2 presents the dataset statistics.
+Table 2 presents the statistics of the dataset.
+Our ELOQ dataset  [^3] https://huggingface.co/datasets/zhiyuanpeng/ELOQ is publicly available.
 3.2. Question Generation
@@
 3).
-For each round:
-  • For all claims in one subset $S_{j}$, we replace them with placeholders “(missing)”.
-  • The modified claims list (with missing claims) is then provided to LLMq without access to $d$. The model is prompted to recover the missing claims based on the remaining claims in the list.
-  • The recovered claims are inserted back into the claims list, replacing the placeholders.
-  • This process is repeated for all subsets $S_{1},…,S_{k}$ across several rounds.
-Finally, we discard claims that are explicitly supported by $d$. This ensures that only truly out-of-scope claims remain in our dataset.
+For each round, we follow steps 3 to 6 in Algorithm 1.
+In practice, we find that even after repeating the hallucination injection process multiple times, some claims remain too general and can still be supported by the document.
+Therefore, we design a prompt-based filter  B.11 to discard these in-scope claims that are explicitly supported by $d$.
+This ensures that only truly out-of-scope claims remain in our dataset.
+Attribute Gold Silver
+# of documents 45 1,955
+# of in-scope questions 103 8,949
+# of out-of-scope questions 113 10,105
+average words per document 280 276
+average words per in-scope question 16 16
+average words per out-of-scope question 16 16
+Table 2. Statistics of ELOQ. Gold is sampled from crawled 2,000 news and annotated by annotators (Section 3.4).
 3.2.3. Question Generation
@@
 In addition to generating out-of-scope questions, we also prompt LLMq to generate in-scope questions for comparison.
-The complete question generation process, along with its prompts, is outlined in Algorithm 1.
-Group Name Cohen’s Kappa Annotator Acc Group’s Agree Acc Ground Truth Acc
+The complete question generation process is outlined in Algorithm 1, and the associated prompts are in Appendix B.
+Group Annotator Cohen’s Kappa Annotator Acc Group’s Agree Acc Ground Truth Acc
 Out-of-Scope Defusion Out-of-Scope Defusion Out-of-Scope Defusion Out-of-Scope Defusion
-1   — 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
-  — 87.50 94.59
-2   — 0.7508 0.8187 93.06 97.14 98.41 96.77
-  — 91.67 90.00
-3   — 0.8615 0.8333 91.67 93.75 98.51 100.00
-  — 98.61 100.00
-Table 3. Agreement scores and the performance of our method on the ELOQ-Gold.
+1 A1 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
+A2 87.50 94.59
+2 A3 0.7508 0.8187 93.06 97.14 98.41 96.77
+A4 91.67 90.00
+3 A5 0.8615 0.8333 91.67 93.75 98.51 100.00
+A6 98.61 100.00
+Table 3.
+Agreement scores and the performance of our method on the ELOQ-Gold.
+“Annotator” represents the annotator’s name.
+“Out-of-Scope” represents whether the question is out-of-scope or not.
+“Defusion” refers to whether an LLM’s response indicates that the question has no suitable answer based on the document.
+Algorithm 2 Evaluation steps
+1:Call LLMr on $(d,q)$ with RAG prompt and get its response $r$ (Appendix B.6,  B.7,  B.8)
+2:Call LLMr $m$ times on $(d,q)$ to check and explain if $q$ is out-of-scope given $d$ (Appendix B.3)
+3:Aggregate $m$ predictions by majority vote
+4:if $q$ is generated as out-of-scope then
+5:     Call LLMq $m$ times on $(d,q,r)$ to check and explain if $r$ defused the confusion (Appendix B.4)
+6:     Aggregate $m$ predictions by majority vote
+7:end if
 3.3. Evaluation
@@
 Evaluation of LLMs’ accuracy in defusing out-of-scope questions from the ELOQ-Silver dataset across diverse news topics.
-For each LLM, the underscored value and stared () value are the minimum and maximum values, respectively, across all the topics within the same prompting method.
+For each LLM, the underscored value and starred () value are the minimum and maximum values, respectively, across all the topics within the same prompting method.
 Bold values are the maximum values for each topic across all the LLMs within the same prompting method.
@@
 4.2. Out-of-Scope Detection
-The main challenge in preventing an LLM from hallucinating or generating incorrect answers to out-of-scope questions is detecting such scenarios.
+The primary challenge in preventing an LLM from hallucinating or generating incorrect answers to out-of-scope questions is detecting such scenarios effectively.
 Once we can accurately identify these cases, we can handle them easily by using a prompt and in-context examples specifically designed for out-of-scope questions.
@@
 4.2.1. Unused-token Classification
-The out-of-scope detection problem can be framed simply as a binary classification task, where we determine if a given query $q$ is out-of-scope or not for a given document $d$.
-The binary label is $y\in\{0,1\}$ where $y=1$ indicates the query is relevant to the document and $y=0$ indicates the query is out-of-scope.
-The objective is to learn a classification function: $f:(q,d)\mapsto p(y=1\mid q,d)$, which estimates the likelihood of the query-document pair to be in-scope.
-We would like to test whether the LLM’s internal features contain enough information to classify a query-document pair effectively to simulate the effect that the LLM “realizes” the out-of-scope scenario before generating a response.
+The out-of-scope detection problem can be framed simply as a binary classification task, where we determine if a given question $q$ is out-of-scope or not for a given document $d$.
+The binary label is $y\in\{0,1\}$ where $y=1$ indicates the question is relevant to the document and $y=0$ indicates the question is out-of-scope.
+The objective is to learn a classification function: $f:(q,d)\mapsto p(y=1\mid q,d)$, which estimates the likelihood of the question-document pair to be in-scope.
+We would like to test whether the LLM’s internal features contain enough information to classify a question-document pair effectively to simulate the effect that the LLM “realizes” the out-of-scope scenario before generating a response.
 To achieve this, we use the sequence representation of the entire input – denoted as $x=\text{concat}(I,q,d,)$ where $I$ is the instruction prompt provided to the LLM (e.g.
@@
 The performance reported in Figure 2 was tested on ELOQ-Gold, the human-labeled data.
-For more details about the training, please refer to Appendix D.
+For more details about the training, please refer to Appendix C.
 We also provide a “Sentence BERT NLI” baseline where we adopt NLI style training using the off-the-shelf  [^4] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 model from HuggingFace.
-Specifically, we follow (Reimers and Gurevych, 2019) to extract document and question [CLS] embeddings $e_{d}$ and $e_{q}$ to use $[e_{q};e_{q};|e_{d}-e_{q}|]$ as the input features.
+Specifically, we follow (Reimers and Gurevych, 2019) to extract document and question [CLS] embeddings $e_{d}$ and $e_{q}$ to use $[e_{d};e_{q};|e_{d}-e_{q}|]$ as the input features.
 4.2.2. Results Discussion
@@
 4.3. Semantic Relevance
-To demonstrate that the generated out-of-scope questions are indeed semantically similar to the document they are generated from, we conduct retrieval experiments using all out-of-scope questions against our document corpus (2000 document).
+To demonstrate that the generated out-of-scope questions are indeed semantically similar to the document from which they are generated, we conduct retrieval experiments using all out-of-scope questions against our document corpus (2,000 documents).
 We employ the following retrieval models: BM25  [^5] https://github.com/castorini/pyserini, SBERT  [^6] https://huggingface.co/sentence-transformers/all-mpnet-base-v2, bge-large-en-v1.5  [^7] https://huggingface.co/BAAI/bge-large-en-v1.5, stella_en_1.5B_v5888https://huggingface.co/NovaSearch/stella_en_1.5B_v5, and Linq-Embed-Mistral [^9] https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral.
@@
 When we relax this condition to Recall@10, around 80% of out-of-scope questions retrieve their associated document within the top 10 results.
-This demonstrates the strong semantic connection between out-of-scope questions and their source documents despite the fact that the document lacks the information needed to answer the question.
+This demonstrates a strong semantic connection between out-of-scope questions and their source documents, despite the fact that the document lacks the necessary information to answer the question.
 Module Accuracy
@@
 Since ELOQ-Gold is a smaller dataset, using $m=9$ did not yield gains, but this doesn’t rule out its potential benefit on the larger ELOQ-Silver dataset.
-Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix C).
-5. Conclusion and Future Works
+Therefore, we adopted $m=9$ in our experiments to balance accuracy and expenses (Appendix A).
+5. Conclusion and Future Work
 We introduced ELOQ, a benchmark for enhancing LLM detection of out-of-scope questions for a given document.
 Human annotation validates that ELOQ achieves 94.91% accuracy in generating out-of-scope and in-scope questions, while our AutoDefusionEval achieves 98.23% accuracy in detecting effective defusion responses.
-We further demonstrated ELOQ’s utility by training out-of-scope detectors.
+We further demonstrated ELOQ’s utility by training out-of-scope detectors, and the results show that a small model (Llama 3.1 8B) can achieve comparable results to the larger one (Llama 3.1 70B) without training.
 For future work, we plan to apply for OpenAI funding to utilize more powerful models, such as o1, to improve both data quality and scale.
-Additionally, we aim to explore post-training techniques to fine-tune LLMs on ELOQ, ensuring they align better with human preferences by clarifying the confusing aspects of a question rather than generating hallucinated answers.
-Appendix A Out-of-scope Example
-Document:
-Justin Jefferson, CeeDee Lamb, NFL Injury Statuses and Fantasy Impact for Week 3 Stephen Maturen/Getty Images
-Two of the best wide receivers in the NFL are expected to be on the field for their respective Week 3 games.
-Justin Jefferson and CeeDee Lamb each received positive prognosis about their injury issues, which is something that can’t be said for the rest of the stars across the NFL.
-Christian McCaffrey is already on injured reserve, Deebo Samuel is out for a weeks and a slew of other running backs and wide receivers are dealing with ailments that could keep them out of Week 3.
-Below is a look at all of the significant injuries that could affect fantasy football matchups across Week 3.
-Justin Jefferson Off Injury Report
-Justin Jefferson was taken off the Minnesota Vikings injury report on Friday.
-Jefferson’s status was up in the air because of a quad injury, but he practiced well enough this week that the injury is not a concern.
-The superstar wide out will be needed for Minnesota’s home clash with the Houston Texans, which has the potential to be a high-scoring affair.
-Jefferson is always the primary target in Minnesota when healthy, but he should have more targets in Week 3 because Jordan Addison and T.J.
-Hockenson are still out.
-Jefferson earned seven targets from Sam Darnold in Week 2.
-Only running back Aaron Jones had more than four targets against the New York Giants.
-Houston’s defense may give Jefferson some fits, led by cornerback Derek Stingley Jr.
-but it has allowed 330 receiving yards to opposing wide outs on just 20 catches through two weeks.
-Jefferson is an automatic start whenever he’s healthy, and his star power may be needed more on certain fantasy rosters in Week 3 depending on how many injuries affect a single roster.
-Question: How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?
-Explanation: This question is out-of-scope as it asks about ”Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft” which was never discussed in the document.
+Additionally, we aim to explore post-training techniques to fine-tune LLMs on ELOQ, ensuring they align better with human preferences by identifying the confusing aspects of a question rather than generating hallucinated answers.
+Acknowledgements.
+We would like to express our sincere gratitude to Xiaoxiao Shang, Ethan Lin, Wei Mo, Zhan Shi, Xinze Ye, and Haisong Wang for their invaluable contributions to data labeling.
+Appendix
+Appendix A Expenses
+Generating the ELOQ dataset using GPT-4o-mini costs approximately \$80.
+For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/.
+Excluding the dataset generation, all other experiments cost around \$100.
 Appendix B Prompt
-B.1. Extract Claims Prompt
+B.1. Extract Claims
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@
 {num_fact} most important facts:
-B.2. Recover Missing Claims Prompt
+B.2. Recover Missing Claims
 Read the document below with a list of {num_fact} facts it contains.
@@
 Completed list of facts:
-B.3. Out-of-scope Judgement Prompt
+B.3. Out-of-scope Judgement
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@
 End your response with ’The answer is: Yes or No.’:
-B.4. Response Defusion Judgement Prompt
+B.4. Response Defusion Judgement
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions. Read the document and answer the question based on the document.
@@
 Answer:
-B.9. In-scope Question Generation Prompt
+B.9. In-scope Question Generation
 You will be provided with a document delimited by triple quotes.
@@
 Questions:
-B.10. Out-of-scope Question Generation Prompt
+B.10. Out-of-scope Question Generation
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@
 Questions:
-B.11. Remove Claims Prompt
+B.11. Remove Claims
 You will be provided with a document delimited by triple quotes. Read the document and follow user’s instructions.
@@
 ”””{document}”””
-{num_true_fact} ground-truth facts
+{num_true_fact} ground-truth facts:
 {true_facts}
@@
 Remaining hallucinated facts:
-Appendix C Expenses
-Generating the ELOQ dataset using GPT-4o-mini costs approximately \$80.
-For LLMs not provided by OpenAI, we used the API service from https://www.together.ai/.
-Excluding the dataset generation, all other experiments cost around \$100.
-Appendix D Embedding Classifier Implementation Details
+Appendix C Embedding Classifier Implementation Details
 We train the classifier on ELOQ-Silver, which is split as: 80% for training, 10% for evaluation, and 10% for testing. After training, we test the classifier on ELOQ-Gold data.
@@
 555 is a random choice.
-The unused-token’s hidden state is used as input features for a 2-layer MLP, which is optimized with a BCE loss.
+The unused tokens ’ hidden state is used as input features for a 2-layer MLP, which is optimized with a binary cross-entropy (BCE) loss.
 A dropout of 0.1 is applied after the first layer.
@@
 The best validation checkpoint is saved and its performance is reported on ELOQ-Gold.
-Acknowledgements.
-We would like to express our sincere gratitude to Xiaoxiao Shang, Ethan Lin, Wei Mo, Zhan Shi, Xinze Ye, and Haisong Wang for their invaluable contributions to data labeling.
+Appendix D Example News
+Justin Jefferson, CeeDee Lamb, NFL Injury Statuses and Fantasy Impact for Week 3 Stephen Maturen/Getty Images
+Two of the best wide receivers in the NFL are expected to be on the field for their respective Week 3 games.
+Justin Jefferson and CeeDee Lamb each received positive prognosis about their injury issues, which is something that can’t be said for the rest of the stars across the NFL.
+Christian McCaffrey is already on injured reserve, Deebo Samuel is out for a weeks and a slew of other running backs and wide receivers are dealing with ailments that could keep them out of Week 3.
+Below is a look at all of the significant injuries that could affect fantasy football matchups across Week 3.
+Justin Jefferson Off Injury Report
+Justin Jefferson was taken off the Minnesota Vikings injury report on Friday.
+Jefferson’s status was up in the air because of a quad injury, but he practiced well enough this week that the injury is not a concern.
+The superstar wide out will be needed for Minnesota’s home clash with the Houston Texans, which has the potential to be a high-scoring affair.
+Jefferson is always the primary target in Minnesota when healthy, but he should have more targets in Week 3 because Jordan Addison and T.J.
+Hockenson are still out.
+Jefferson earned seven targets from Sam Darnold in Week 2.
+Only running back Aaron Jones had more than four targets against the New York Giants.
+Houston’s defense may give Jefferson some fits, led by cornerback Derek Stingley Jr.
+but it has allowed 330 receiving yards to opposing wide outs on just 20 catches through two weeks.
+Jefferson is an automatic start whenever he’s healthy, and his star power may be needed more on certain fantasy rosters in Week 3 depending on how many injuries affect a single roster.